### v1

# Chatbot Companionship: A Mixed-Methods Study of Companion Chatbot Usage Patterns and Their Relationship to Loneliness in Active Users

Abstract
As artificial intelligence becomes increasingly sophisticated, companion chatbots have been proposed as a potential solution to the growing epidemic of loneliness. However, the impact of these AI companions on users’ psychological well-being and social behaviors remains poorly understood. This study presents a large-scale survey (n = 404) of regular users of companion chatbots, investigating the relationship between chatbot usage and their experience of loneliness. We find a small but significant direct correlation between session length with chatbots and loneliness, and we develop a model of this relationship through multiple regression analysis, finding social attraction and neuroticism as moderators. We find seven clusters of users, including socially fulfilled dependent users to lonely dependent users. Our work contributes to the ongoing dialogue about the role of AI in social and emotional support, offering insights for what kind of human-AI connections might lead to emotional well-being and complement rather than replace human connections.
Keywords Companion chatbots, loneliness, artificial intelligence, artificial agents, human-AI interaction, social support
Figure 1: Overview of the study’s methodology and research questions. A survey of 404 participants explored chatbot usage and its psychological impacts. RQ1 investigates the motivations and uses of chatbots through descriptive statistics and free responses. RQ2 examines the relationship between chatbot usage and loneliness using correlation and multiple regression analyses, identifying potential mediating and moderating factors. RQ3 focuses on identifying distinct user profiles through K-means cluster analysis, considering psychological and social factors. The results highlight the complexity of human-AI interactions and their varied impacts on emotional well-being.
1 Introduction
In an increasingly digital world, we experience both increased connectivity and increased social isolation. Despite unprecedented technological means to connect, loneliness remains a pervasive issue, affecting approximately a third of individuals in industrialized countries [Loveys et al.(2019)]. This phenomenon has led to what the U.S. Surgeon General declared an “epidemic of loneliness” in May 2023 [Office of the Assistant Secretary for Health (OASH)(2023)]. The gravity of this issue is underscored by research indicating that chronic loneliness can increase mortality risk by 26-29%, comparable to smoking 15 cigarettes a day or having an alcohol use disorder [Chawla et al.(2021)].
Artificial intelligence (AI) has emerged as a potential tool for addressing loneliness. Companion chatbots, chatbots engage in complex conversations and maintain persistent relationships with users, have gained significant popularity, with platforms like Replika and Character.ai reporting over 30 million and 20 million users, respectively [Patel(2024), Shewale(2024)]. These AI companions offer users the opportunity for constant, judgment-free interaction. However, the impact of these artificial relationships on users’ psychological well-being and social behaviors needs further study.
Some research suggests that AI companions, even when used over the short term, decreases feelings of loneliness [Loveys et al.(2019), Gasteiger et al.(2021), De Freitas et al.(2024)]. However, concerns have been raised about the authenticity of these relationships, suggesting that they lower the bar for what people think relationships can be [Turkle(2011)]. This could lead to unrealistic expectations and further social withdrawal.
The nuanced nature of human-AI interaction in this context presents a critical gap in our understanding. Specifically, we lack comprehensive knowledge about the factors that determine whether and to what extent companion chatbots impact loneliness and other psychological and social factors. As Xia et al. [Xia et al.(2024)] noted, artificial agents may complement human social interactions in some contexts while replacing them in others. This variability in outcomes underscores the need for a more comprehensive and granular analysis of user characteristics, usage patterns, and their relationships to psychological well-being.
To address this gap, we conducted a large-scale survey (n = 404) of regular companion chatbot users, with the following research questions:
  • RQ1: What motivates the use of companion chatbots, and what do people use them for?
  • RQ2: What is the relationship between companion chatbot usage and loneliness, and what factors mediate or moderate this relationship?
  • RQ3: Can we identify and characterize distinct user profiles based on usage patterns, loneliness levels, problematic use, and other psychological and social factors?
Our study (see Figure 1) aims to provide a more nuanced understanding of the impact of companion chatbots, moving beyond binary assessments of positive or negative effects. By identifying the characteristics of users who experience psychosocial benefits from chatbot interactions (such as reduced loneliness, improved social skills, and better mental well-being) and those who may experience negative consequences (such as emotional dependence and social withdrawal), we hope to inform more targeted and ethical design practices. For example, more information could guide the development of customizable interfaces that emphasize features beneficial to specific user groups or implement safeguards for at-risk users. By examining how users’ mental models and expectations of AI companions influence outcomes, we can explore strategies for fostering healthy human-AI relationships that complement rather than replace human connections.
The ethical implications of widespread AI companion use are multifaceted. While these technologies may serve as a valuable intervention for individuals with limited access to social support, there are concerns about the long-term effects on social skills and expectations for human relationships. Our study aims to provide empirical grounding for these discussions, informing both design practices and policy considerations in the rapidly evolving landscape of AI-mediated social interaction.
Through a combination of statistical analyses, including multiple regression and cluster analysis, along with qualitative assessment of user responses, we present the following contributions:
  • A model for understanding the relationship between companion chatbot usage and loneliness, accounting for mediators such as socialization and moderators of social attraction and neuroticism.
  • A typology of companion chatbot users that takes into account usage patterns, psychological characteristics, and perceived benefits or drawbacks.
  • An exploration of motivations and usage characteristics of users of companion chatbots.
  • Implications for creating more effective and ethically responsible AI companions that enhance well-being.
By addressing these questions, we aim to advance the field of human-computer interaction in the domain of AI companionship, providing insights that can guide the development of technologies that genuinely enhance human well-being and social connection in our increasingly digital world.
2 Background and Related Work
2.1 Loneliness and Social Support
Loneliness, defined as the subjective feeling of isolation regardless of one’s actual social network, has become a pressing global issue affecting approximately a third of individuals in industrialized countries [Loveys et al.(2019)]. The U.S. Surgeon General’s declaration of loneliness as an epidemic in 2023 underscores the gravity of this issue [Office of the Assistant Secretary for Health (OASH)(2023)]. Chronic loneliness is associated with a significantly increased risk of mortality—approximately 26-29%, comparable to smoking 15 cigarettes a day or having an alcohol use disorder [Freedman and Nicolle(2020)].
Various risk factors contribute to social isolation and loneliness, including socio-demographic factors (e.g., increased age, low income), medical factors (e.g., cognitive impairment, sensory loss), and social factors (e.g., living alone, lack of transportation) [Freedman and Nicolle(2020)]. Older populations are particularly vulnerable due to the accumulation of these risk factors and potentially lower digital literacy, which can limit their use of modern communication technologies [Gasteiger et al.(2021)].
Social support has been identified as a strong predictor of good physical and mental health, as well as resilience to stress. It can be defined as the support accessible to an individual via their social ties to others including individuals, groups, and a larger community [Ozbay et al.(2007)]. Importantly, the quality of relationships is a better predictor of health outcomes than the mere quantity of social connections [Wang et al.(2018), Schutter et al.(2021)].
2.2 Loneliness Coping Strategies and Interventions
Individuals employ various strategies to cope with loneliness, ranging from traditional approaches to digital solutions. Rokach (1998) identified six primary factors in coping with loneliness: Reflection and Acceptance (involving solitary reflection and acceptance of loneliness), Self-Development and Understanding (involving increased understanding of oneself through relationships and therapy), Social Support Network (increased social involvement), Distancing and Denial (involving unhealthy behaviors such as turning to drugs and self-harming), Religion and Faith (involving the feeling of belonging in religious communities), and Increased Activity (devoting more time to extracurricular activities) [Rokach and Brock(1998)]. A systematic review by Deckx et al. (2018) further categorized coping strategies into problem-focused and emotion-focused approaches, with problem-focused strategies consistently associated with lower levels of loneliness [Deckx et al.(2018)].
In recent years, digital platforms have emerged as commonly used coping mechanisms for loneliness, providing both social interaction and passive engagement. Individuals may use video platforms such as YouTube to cope with their loneliness – higher loneliness is associated with the formation of parasocial relationships, where individuals form one-sided relationships with content creators [Niu et al.(2021), de Bérail et al.(2019), Balcombe and Leo(2023)]. This is consistent with the parasocial compensation hypothesis, which postulates that lonely, socially isolated, and socially anxious individuals tend to be more engaged in parasocial relationships [Madison et al.(2016)].
Online communities and forums, such as Reddit, Facebook Groups, and Twitch, facilitate real-time communication and foster social connections across shared interests, allowing individuals to engage in virtual communities and potentially alleviate social isolation [Seering et al.(2019)]. Existing work suggests that seeking and offering social support are key activities in forums [Johnson et al.(2020)]. However, high levels of loneliness and social anxiety are associated with addiction and problematic use of these platforms, mediated by increased social comparison and fear of missing out [Reer et al.(2019), O’Day and Heimberg(2021), Burke et al.(2010)]. Interestingly, image-based media like Instagram may decrease loneliness due to the enhanced intimacy of image-sharing, while text-based media such as X (Twitter) does not show the same effect [Pittman and Reich(2016)].
There are, however, effective interventions for reducing loneliness. These typically fall into four categories, and can be delivered in-person or online: (1) modifying maladaptive social cognition (e.g., cognitive behavioral therapy, psychodynamic therapy, reminiscence, and mindfulness), (2) improving social support (e.g., involving care professionals in health and social care provision), (3) facilitating social interaction (e.g., clubs, shared interest groups, and videoconferencing), and (4) enhancing social skills (e.g., training programs and computer courses) [Rokach and Brock(1998), Gasteiger et al.(2021), Hickin et al.(2021)]. While these interventions can reduce loneliness, their efficacy varies across individuals. Hickin et al. [Hickin et al.(2021)] conducted a systematic review and meta-analysis of psychological interventions for loneliness, finding that such interventions are generally effective across the lifespan. However, they emphasize that the considerable heterogeneity in intervention effectiveness underscores the importance of personalized approaches. This finding aligns with the growing recognition that lonely individuals are a diverse group with varying needs [Cacioppo and Cacioppo(2018)], suggesting that interventions – including technological ones – must be tailored to individual circumstances for optimal impact.
2.3 Companion Chatbots
The emergence of companion chatbots, also known as virtual friends or digital companions, represents a novel approach to tackling loneliness [Wang et al.(2024), Goodings et al.(2024), Boine(2023), Chen et al.(2024), Shani et al.(2022), Brewer(2022)]. We refer to these as commercially available chatbots that employ generative AI to engage in complex conversations and that maintain persistent relationships with users. These companion chatbots are accessible to a broad demographic, with several popular companion chatbot services (see Figure 2) having gained significant user bases:
Figure 2: Examples of companion chatbots: Replika, Character.ai, Snapchat’s My AI, Anima, and Pi.
  • Replika: An app featuring a generative AI chatbot designed to act as a friend, sibling, assistant, or romantic partner. Users interact with one chatbot that adaptively develops over time through continued conversations. Replika’s CEO reported over 30 million registered users worldwide in August 2024 [Patel(2024)], with its audience primarily aged 25-30 [Christopher(2023)].
  • Character.ai: Established in late 2022, this platform allows users to role-play using constructed personas and interact with chatbots based on fictional characters. With over 20 million users reported in March 2024, its audience primarily comprises young adults aged 18-24 [Shewale(2024)]. Notably, the sixth most popular character on the platform is “Psychologist,” with over 70 million messages shared in a year, indicating a significant trend of users seeking mental health support from AI [Tidy(2024)].
  • Snapchat’s My AI: A chatbot integrated into the Snapchat platform, powered by large language models like GPT-3.5 and GPT-4. It offers advice, answers questions, and assists with various tasks, demonstrating the integration of AI companions into mainstream social media platforms. Over 150 million individuals have used My AI since its launch [Hutchinson(2023)].
  • Anima: A chatbot service that enables role-playing conversations with different personas, allowing users to explore various scenarios and characters.
  • Pi: Developed by Inflection AI, Pi is an AI companion designed to engage in open-ended conversations on a wide range of topics. It aims to be a supportive, empathetic presence for users, offering advice, emotional support, and intellectual discussions, with six million monthly active users reported in March 2024 [Wodecki(2024)].
The popularity of these services, particularly among young adults, speaks to their growing acceptance as sources of social interaction and emotional support [Koulouri et al.(2022), Xygkou et al.(2024), Xygkou et al.(2023)]. Research suggests that artificial companions can reduce loneliness through various mechanisms, including direct companionship, acting as a catalyst for social interaction, facilitating remote communication, and providing reminders for social activities [Gasteiger et al.(2021)].
The perceived “realness” of the connection with AI companions is a crucial factor in their effectiveness. Users who perceive chatbots as having human-like qualities, such as consciousness or agency, often experience improved social health and reduced loneliness [Xia et al.(2024)]. This anthropomorphism can enhance the effectiveness of companion chatbots in mitigating feelings of isolation, though it carries a risk of leading to the user dehumanizing other people [Zehnder et al.(2021)]. Pataranutaporn & Liu et al. [Pataranutaporn et al.(2023)] also found that mental models altered experience – priming users with different beliefs about an AI’s motives (caring, manipulative, or neutral) significantly affected their assessment of the AI’s trustworthiness, empathy, and effectiveness. This suggests that users’ mental models of AI systems play a crucial role in shaping their interactions and experiences.
Previous work has suggested that the ability to empathize is inversely correlated with loneliness over a lifetime [Beadle et al.(2012)], but the question remains of what can be the role of a chatbot that simply performs empathy. Role-play has been shown to increase empathy [Bearman et al.(2015), Dewi et al.(2019)], and the potential of AI-facilitated role-play to influence empathy and, by extension, loneliness, represents an unexplored area in the research literature, despite the large number of users engaging in such interactions.
Existing research on the influence of artificial agents on socialization with other people has produced mixed results, with the work mostly qualitative in nature [Van der Loos(2014), Turkle(2011)]. One study suggested that artificial agents may end up replacing humans for task-oriented interactions, while in social contexts they are more likely to complement human social interactions [Xia et al.(2024)].
2.4 Human-AI Interaction for Emotional Support
Recent studies suggest that even limited amounts of AI companionship reduce the experience of loneliness [De Freitas et al.(2024)]. However, the long-term effects and potential risks of relying on AI for emotional support remain subjects of ongoing investigation and ethical debate [Li et al.(2023), Abd-Alrazaq et al.(2020)].
Certain populations, such as LGBTQ individuals who may lack support from family and friends, are increasingly turning to LLM-based chatbots for mental health support. These users often seek out chatbots for immediately accessible support, viewing conversations with AI as safe spaces for emotional expression, showing the potential for AI to provide a sense of connection and safety for marginalized groups [Ma et al.(2024)].
Recent studies by Jo et al. [Jo et al.(2023), Jo et al.(2024)] explored the deployment of an LLM-driven chatbot, CareCall, for public health interventions targeting socially isolated individuals. Their findings suggest that chatbots can alleviate public health workloads and mitigate loneliness, with users perceiving the system as personal and emotionally supportive. However, challenges such as trust, privacy, and the limitations of LLM-driven chatbots in addressing complex health needs were also highlighted in this study, emphasizing the importance of careful system design and managing stakeholder expectations.
Neves et al. [Neves et al.(2015)] examined the difficulties faced by older adults in adopting new communication technologies, highlighting the importance of user-centered design and support systems in facilitating technology acceptance. These insights are crucial when considering the potential of AI chatbots as a scalable solution for emotional support and loneliness reduction.
While companion chatbots offer promising avenues for emotional support, there are significant ethical considerations and potential risks associated with their use. These include the possibility of over-reliance on AI, consequences for mood and affect, delays in seeking professional help, and further withdrawal from human socialization [Pataranutaporn et al.(2021)]. The development of dependence on AI agents and the potential replacement of real human relationships raise important questions about the long-term societal impacts of widespread AI companion use.
The tensions between the benefits and risks of AI companionship underscore the need for a nuanced understanding of how different users interact with and are affected by these technologies. As AI continues to advance, it becomes increasingly important to investigate the factors that determine positive or negative outcomes from companion chatbot use. This aligns with the principles of positive computing in human-AI interaction, which aims to design technology that supports psychological well-being [Calvo and Peters(2017)]. Towards this goal, our research seeks to inform the design of more effective and ethically responsible AI companions that support rather than supplant human social connections.
More specific, granular information about users and their needs is needed. Our research seeks to identify the characteristics of users who benefit from chatbot interactions, as well as those who may experience negative consequences. Our intent is to inform more targeted and ethical design practices in the rapidly evolving landscape of human-AI interaction.
3 Methods
We surveyed companion chatbot users recruited from CloudResearch, administering several psychological scales and measures of chatbot usage frequency and session length. Using the Qualtrics sample size estimator, we roughly estimated a minimum size of 385 with a 95% confidence level and a 5% margin of error. Participants were first recruited with a screening survey in which they were asked about whether or not they use companion chatbots regularly (at least once a week for at least a month). They were asked to elaborate on their experiences with chatbots briefly. Participants could only access the screening survey on CloudResearch if they were fluent in English and had prior experience using AI. A researcher manually read through the respondents’ brief free responses to assess whether they were engaging with companion chatbots regularly for social and emotional purposes. Those who were determined to be regular users were approved and given access to the full survey. Participants were screened and recruited through multiple rounds until we reached our target pool of qualified individuals. All qualifying participants were then invited to complete the main survey. We intentionally over-recruited to account for incomplete responses. After filtering out incomplete responses from the main survey, the final sample consisted of 404 valid responses (n = 404), slightly exceeding our initial target of 400. This study was approved by the institution’s IRB.
We conducted the following analyses:
  • Participant Demographics: We calculated descriptive statistics to create a general overview of our participants.
  • Correlation Analysis: We calculated a Spearman correlation matrix with Benjamini-Hochberg correction as a starting point for exploring relationships in our numerical variables.
  • Multiple Regression Analysis: We performed an exploratory multiple regression analysis to construct a more nuanced model that can capture relationships between variables not reflected in the correlation analysis.
  • Participant Cluster Analysis: We performed K-means clustering of our numerical data to find distinct groups of users in our data.
  • Chatbot Usage Characteristics: We calculated descriptive statistics on multiple-choice questions regarding companion chatbot usage and examined participants’ free responses, both overall and grouped by the clusters we found. Our intention was to characterize our participants overall and provide more specific characteristics and examples of responses for each cluster.
These analytical approaches were designed to address our primary research questions:
  1. 1.
    To explore motivations and uses of companion chatbots (RQ1), we analyzed the descriptive statistics of usage characteristics and examined participants’ free responses.
  2. 2.
    To investigate the relationship between chatbot usage and loneliness (RQ2), we employed correlation analysis and multiple regression analysis. These methods allowed us to identify potential mediating or moderating factors in this relationship.
  3. 3.
    To identify and characterize distinct user profiles (RQ3), we conducted K-means cluster analysis on our numerical data, including usage patterns, loneliness levels, and other psychological and social factors. We then examined chatbot usage characteristics within each cluster to provide more nuanced profiles.
Figure 3: Comprehensive flowchart detailing the multi-method approach employed in this study. The diagram outlines the sequence of analytical steps, including participant recruitment, survey administration, data processing, correlation analysis, multiple regression modeling, cluster analysis, and the interpretation of both quantitative results and free-text responses.
We summarize these methods in Figure 3. By combining these quantitative and qualitative approaches, we aimed to provide a comprehensive understanding of companion chatbot usage patterns, their relationship with psychological factors such as loneliness, and the characteristics of different user groups.
3.1 Survey Design
Our survey instrument was designed to investigate the relationship between chatbot usage and loneliness, as well as potential mediating factors. The survey comprised both established scales (some adapted) and custom questions developed for this study. The full survey can be found as a Word document in Supplementary Material.
3.1.1 Established Scales and Adaptations
We included several established scales, many of which were adapted to better fit our study context and to manage survey length:
  • UCLA Loneliness Scale (ULS-8) [Hays and DiMatteo(1987)]: Measures loneliness.
  • Lubben Social Network Scale (LSNS-6) [Lubben et al.(2006)]: Measures social isolation.
  • Multidimensional Scale of Perceived Social Support (MSPSS) [Zimet et al.(1990)]: Assesses perceived social support.
  • Brief Rosenberg Self-Esteem Scale (B-RSES) [Monteiro et al.(2022)]: Evaluates self-esteem.
  • Big Five Inventory (BFI-10) [Rammstedt and John(2007)]: Evaluates personality traits based on five factors: extraversion, neuroticism (tendency to towards anxiety and other negative feelings), agreeableness (tendency to be compassionate and cooperative), conscientiousness (tendency to be self-disciplined and goal-directed), and openness to experience (tendency to be open-minded, imaginative, and insightful) [McCrae and Costa(1999)].
  • Attitudes Towards AI Scale [Schepman and Rodway(2020)]: Measures attitudes towards artificial intelligence. Used the adaptation found in [Pataranutaporn et al.(2023)].
  • Human-Chatbot Interaction Effect Scale [Xia et al.(2024)]: Measure of the effect of human-chatbot interaction on human-human interaction. Measures the respondent’s extraversion, expectations for chatbot services (in performance and meeting their needs), trust in interactions with social chatbots, perceived value of substituting people with chatbots, trust in interactions with people, and interpersonal communication intention (motivation and willingness to interact with people). Changed phrasing to refer to only chatbots and people rather than service providers. Removed the extraversion subscale, shortened subscales based on factor loadings, and made two versions of the interpersonal communication intention subscales, one for familiar people and one for unfamiliar people.
  • Generalized Problematic Internet Use Scale (GPIUS-2) [Caplan(2010)]: Shortened to 6 items based on relevance and subscale representation. Rephrased to reference chatbot use instead of general internet use.
  • State Empathy Scale [Shen(2010)]: Adjusted wording to be more general for application to both chatbots and people.
  • Interpersonal Attraction Scale [McCroskey and McCain(1974)]: A scale measuring the respondent’s social attraction (desire to have a friendship and spend time with another), physical attraction (attraction to appearance), and task attraction (inclination to work with another based on competence and dependability). Shortened each of the three subscales to four items based on factor loadings and made two versions, one for chatbots and one for a close person; the version for a close person removed the physical attraction subscale due to the potential of the respondent’s close person being a family member.
  • Perceived Homophily Scale [Mccroskey et al.(1975)]: A measurement of perceived similarity in interpersonal communication, which is associated with increased communication effectiveness. Changed to unidirectional Likert statements (e.g., “like me” instead of “unlike me – like me”). Selected 6 items from potential items based on factor loadings. Rephrased slightly into phrases and adjusted phrasing to make two versions, one for chatbots and one for a close person.
  • Attributional Confidence Scale [Gudykunst and Nishida(1986)]: A measure of the respondent’s ability to predict others’ attitudes, feelings, and behavior [Clatterbuck(1979)]. Shortened to 5 items, removing items irrelevant to chatbots.
  • System Usability Scale (SUS) [Brooke(1996)]: A measurement of the respondent’s subjective assessment of a system’s usability. Adjusted phrasing to refer specifically to chatbots.
These scales were selected to provide a comprehensive view of participants’ social and psychological characteristics that might mediate the relationship between chatbot usage and loneliness. The adaptations were made to ensure relevance to the chatbot context and to manage survey length, aiming to reduce participant fatigue given the comprehensive nature of our investigation.
3.1.2 Custom Survey Items
In addition to established scales, we developed custom items to capture specific aspects of chatbot usage:
  • Usage Characteristics: Frequency of chatbot use and average session length.
  • Motivations and Continued Use: Multiple-choice questions with required free-text elaboration on why participants began using chatbots and why they continue to use them.
  • Human vs. Chatbot Preference: Several items exploring participants’ preferences for human or chatbot interaction in various contexts.
  • Conversation Topics: Questions about the topics participants discuss with chatbots.
  • Perceived Impact on Relationships: Likert-scale items assessing participants’ perceptions of how chatbot use has affected their relationships with people.
These custom items were designed to provide context-specific data on chatbot usage patterns and perceived effects, complementing the insights gained from established scales.
3.2 Correlation and Multiple Regression
Data processing and statistical analyses were performed using Python, with the statsmodels library for regression analyses and scikit-learn for data preprocessing. We used a significance level of $\alpha=0.05$ for all statistical tests, unless otherwise specified. Respondents’ frequency of chatbot usage and session length were converted to linear numerical scores; any multiple responses were averaged. Numerical variables, including the frequency and session length score, were scaled to have a mean of 0 and a standard deviation of 1 using StandardScaler to improve interpretability and reduce potential multicollinearity [Aiken and West(1991)]. We focused on chatbot session length as the independent variable representing chatbot usage in the model, as it had a significant correlation with loneliness in the correlation analysis where frequency did not. We computed a Spearman correlation matrix for all standardized variables and calculated corresponding p-values to identify significant relationships (p < 0.05) between variables of interest [Cohen et al.(2003)]. The p-values were then corrected using the Benjamini-Hochberg procedure [Benjamini and Hochberg(1995)]. For the following analyses, we focused on loneliness (ULS-8) as the dependent variable and session length score as the primary independent variable, though frequency score was also included in our process of model exploration.
Potential mediators were identified using the corrected p-values (p < 0.05). We conducted mediation analysis following Baron and Kenny’s [Baron and Kenny(1986)] approach, regressing the dependent variable on the independent variable, the mediator on the independent variable, and the dependent variable on both the independent variable and mediator. We calculated indirect effects and conducted Sobel tests [Sobel(1982)] to assess the significance of mediation (p < 0.05).
We identified potential moderators using the corrected p-values (p < 0.05). To assess moderation effects, we conducted moderation analyses following the approach outlined by Aiken and West [Aiken and West(1991)]. For each potential moderator, we created an interaction term with the independent variable and fitted a linear regression model including the independent variable, moderator, and their interaction. The significance of the interaction term (p < 0.05) was used to determine the presence of moderation effects.
To account for potential confounding variables, we implemented a data-driven approach to identify and control for confounders [Greenland et al.(1999), VanderWeele and Shpitser(2013)]. We examined the correlation matrix and corresponding p-values to identify variables that were significantly correlated (corrected $p<0.05$) with both the independent and dependent variables.
To address potential multicollinearity, we calculated Variance Inflation Factors (VIF), considering variables with VIF > 5 for elimination [Hair et al.(2010), O’brien(2007)]. After removing the state empathy subscales, which were redundant with the full scale, all predictors were below the threshold.
We explored various regression models using Ordinary Least Squares (OLS) regression, beginning with all identified mediators, moderators, and confounders. We included socialization (LSNS) as a potential mediator and neuroticism as a potential confounder on a theoretical basis, as socialization is the objective counterpart to loneliness, and neuroticism is strongly linked to loneliness [Buecker et al.(2020)] and may act as a driver for emotional regulation using chatbots. While retaining age and sex as control variables, we iteratively refined the model by removing predictors with high p-values (generally p > 0.1) in a process of backward elimination, to make a more parsimonious and interpretable model [Cohen et al.(2003)]. After elimination, we experimented with adding social attraction to a close person, and its interaction with session length, to the model. For each of the remaining predictors, we explored adding one new interaction term at a time, with significant interaction becoming part of the model. Frequency score was removed from the model, as its p-value (0.636) was far above the significant threshold. Categorical variables (area type - rural/urban/suburban, education, relationship/marital status, political party, race, employment status) were tested using ANOVA one at a time but were not retained in the final model due to non-significance.
We assessed model fit using R-squared, F-statistic, and condition number. Model assumptions were checked using Q-Q plots, residuals vs. fitted plots, and influence plots [Cook and Weisberg(1982), Fox(2015)]. The summary of the final model from our exploration can be seen in Figure 14.
To examine the potential mediators of our final model, we conducted a parallel multiple mediation analysis. This analysis was performed using a bootstrapping approach with 5,000 resamples, as recommended by [Preacher and Hayes(2008)]. The analysis controlled for several covariates, including age, sex, and personality traits (BFI). Confidence intervals (95%) for the indirect effects were calculated using the percentile method. All analyses were conducted using Python with the statsmodels library.
3.3 Cluster Analysis
To identify distinct user profiles of chatbot users based on their psychosocial characteristics, we employed K-means clustering analysis. These profiles were identified based on a set of features with significant correlations (p > 0.05, corr > 0.3) to key variables, as well as features relevant to chatbot usage including usage frequency and measures of empathy and attraction with a chatbot; the features were reduced to 10 principal components. After some exploration, we used Chatbot Session Length, Frequency, Loneliness (ULS-8), and Problematic Use of Chatbot (GPIUS-2) for our key variables. The optimal number of clusters was determined using a combination of quantitative metrics and qualitative interpretation:
  1. 1.
    We calculated the Within-Cluster Sum of Squares (WCSS) for a range of cluster numbers (2-14) and plotted the results to visualize the elbow method [Kodinariya and Makwana(2013)].
  2. 2.
    We computed silhouette scores for the same range of clusters to assess cluster cohesion and separation [Kaufman and Rousseeuw(2009)].
  3. 3.
    We examined the interpretability and meaningfulness of the resulting clusters for different k values, considering our research questions and domain knowledge.
The silhouette scores increased slightly as the number of clusters increased from 3 to 7, dropping to a local minimum at 8 clusters. The silhouette score fluctuated without consistent increases and there was no clear bend in the elbow method, we ultimately chose a 7-cluster solution. This decision was also based on the interpretability and distinctiveness of the resulting user profiles, aligning with our objective to understand the characteristics of different groups of users.
The final clustering was performed using the K-means algorithm with k=7, implemented in Python using the scikit-learn library [Pedregosa et al.(2012)]. The resulting clusters were then characterized based on their centroids and distribution of key variables.
To validate our clustering solution, we employed a multi-faceted approach:
  1. 1.
    We conducted Kruskal-Wallis tests to examine differences among clusters for key variables (Session Length, Loneliness, and Problematic Use), as the data violated normality assumptions [Kruskal and Wallis(1952)].
  2. 2.
    We calculated the average silhouette score to assess the quality of cluster separation [Rousseeuw(1987)].
  3. 3.
    We performed a Multivariate Analysis of Variance (MANOVA) to test for overall differences among clusters across multiple dependent variables simultaneously [Huberty and Olejnik(2007)].
  4. 4.
    We used Box’s M-test to check the assumption of homogeneity of covariance matrices for the MANOVA [Box(1949)].
These validation techniques provide a comprehensive assessment of the cluster solution’s statistical robustness and meaningful differentiation among user profiles.
3.4 Interpretation of Quantitative Results and Free-Text Responses
To provide a comprehensive understanding of our findings, we employed a two-pronged approach to interpreting the results of our quantitative analyses. First, we calculated descriptive statistics on companion chatbot usage characteristics for both our entire dataset and the clusters we identified through K-means analysis. This provided a quantitative overview of usage patterns and user profiles.
Second, we examined the free-text responses associated with multiple-choice items to complement our quantitative findings. While not a formal qualitative analysis, this examination allowed us to contextualize our quantitative results and provide illustrative examples of user experiences within each cluster and across the entire sample.
By integrating these approaches, we were able to capture both broad trends identified through our quantitative analyses and individual nuances in user experiences expressed in their own words. This method provided a richer interpretation of the data, offering insights into the motivations, perceptions, and experiences of companion chatbot users that might not be fully captured by quantitative measures alone.
4 Results
4.1 Participant Demographics and Usage Characteristics
We surveyed 404 companion chatbot users recruited via CloudResearch. Participants ranged in age from 18 to 73, with the distribution shown in Figure 5, and were primarily from the United States. Additional demographic characteristics are visualized in Figure 4.
In addition to the participants’ demographics, we also collected information about their usage, including which companion chatbots they use, how much they use the chatbots, why they use the chatbots, and what they use the chatbots for. Most participants utilize Snapchat MyAI (31.63%), Character.ai (30.39%), or Replika (22.61%) (see Figure 6). Most respondents reported spending 15 and 30 minutes in each session with a companion chatbot (42.64%), with 35.39% reported sessions shorter than 15 minutes. A smaller portion of users engage for 30 minutes to an hour (15.14%), and only 4.90% of participants reported sessions lasting between 1 and 2 hours. Very few users spend extended periods chatting, with only 1.28% of sessions exceeding 4 hours, and 0.64% lasting between 2 and 4 hours. Additional characteristics can be seen in Figure 6 and Figure 7.
Figure 4: Comprehensive demographic breakdown of study participants. Six pie charts illustrate the distribution of gender, race, relationship/marital status, household income, country of residence, and employment status.
Figure 5: Participants’ age distribution. Participants were 18 to 73 years of age.
4.1.1 Motivations for initial interest and continued use of companion chatbots.
The reasons participants use and continue to engage with companion chatbots are diverse (see Figure 7), with non-social motivations predominating. Curiosity about AI and chatbots (30.89% initial use, 20.91% continued use) is a primary driver. As one user stated, “I was curious to see what they were like,” capturing the initial intrigue many feel. Others sought to explore the technology’s capabilities: “I wanted to see if it could actually emulate a real person.”
Figure 6: Overview of chatbot usage patterns among participants. The graph on the left shows the most used chatbot platforms, with 31.6% of participants using Snapchat MyAI, 30.4% using Character.ai, and 22.6% using Replika, while smaller portions use other platforms such as Anima (4.8%) and Xiaoice (1.2%). The center graph presents the chatbot session length, where 42.6% of participants use chatbots for 15–30 minutes, 35.4% for less than 15 minutes, and 15.1% for 30 minutes to 1 hour. The right graph displays chatbot usage frequency, with 43.9% using chatbots for a few months, 35.2% using them for several months to around a year, and smaller percentages for longer durations such as multiple years (7.7%).
Figure 7: Analysis of participant responses regarding their use of companion chatbots. The top-left graph presents reasons for using companion chatbots, where 30.9% of participants expressed curiosity about AI and chatbots, followed by 26.4% who sought fun and entertainment. The top-right graph displays conversation topics with chatbots, with casual conversation being the most common (26.3%), followed by entertainment and play (21.7%). The bottom-left graph shows motivations for continued use, where enjoyment/fun was the leading factor (23.7%), followed by satisfying curiosity (20.9%). Lastly, the bottom-right graph illustrates preferred devices for companion chatbot use, with the majority of users favoring smartphones (36.9%), followed by computers (30.9%).
Fun and entertainment (26.35% initial, 23.69% continued) are significant factors. Users enjoy role-playing with characters and find humor in interactions: “I’ve watched YouTubers use Character.ai before and found it to be a humorous experience.” The entertainment often intertwines with creativity (17.81% initial use), with users employing chatbots for storytelling and game preparation: “I use chatbots to practice conversations for my D&D game, creating NPC personalities and dialogue.”
Passing time and reducing stress (12.88% continued use) are notable reasons for ongoing engagement. Users find chatbots helpful for relaxation: “When I’m stressed or bored, chatting with an AI feels like a safe and fun way to distract myself.”
Social motivations, while less prevalent, are significant for some users. These include coping with loneliness (12.24% initial, 6.92% continued), improving social and conversational skills (7.24% initial), and seeking companionship (9.86% continued). One user noted, “I started using Replika during the COVID-19 pandemic when I was lonely and seeking some form of social interaction.” For others, chatbots offer a low-pressure environment to practice social skills: “I’m autistic, so I’m not the best at social interaction, and I found that practicing with chatbots helped.”
A small portion of users (3.53% initial) seek relationship-like interactions with chatbots, exploring emotional connections in a safe space. Continued use is also motivated by emotional experiences (6.52%), self-discovery (6.12%), understanding others (5.88%), and boosting self-confidence (5.41%). One user shared, “I’ve learned a lot about myself from talking to the chatbot – it helps me think through things in a different way;” another noted that “Chatbots give me different perspectives that I wouldn’t have considered on my own, which helps me when interacting with others in real life,” demonstrating how the social simulation offered by companion chatbots offer the opportunity to improve both interpersonal and intrapersonal understanding.
4.1.2 Topics of conversation with companion chatbots
The range of topics discussed with chatbots is diverse, reflecting users’ varied interests and needs. Casual conversations (26.28%) dominate, with users engaging in everyday chat without social pressure: “I often chat with my AI about random things like my day, and it feels like I’m having a normal conversation without the pressure of talking to a real person.”
Entertainment and play (21.66%) are common, especially for users interested in role-playing or creative writing: “I use chatbots for world-building inspiration—creating characters and scenarios for my Dungeons and Dragons campaigns.”
Personal issues and mental health (14.17%) are notable topics, with users valuing the non-judgmental nature of chatbots: “I can talk about my problems, and it’s like having a private conversation with no fear of being criticized.” One user shared how a companion chatbot becomes an option for venting when people are not an option, stating, “I struggle with social anxiety, and when I can’t talk to anyone else, I use a chatbot to vent my feelings.”
Users also discuss past events (12.60%) and future plans (11.70%), using chatbots as tools for reflection and planning. One user explained, “I talk to the chatbot about my future goals and plans because it helps me organize my thoughts.”
Interpersonal issues and drama (10.96%) are explored with chatbots, allowing users to gain perspective on social conflicts: “When I have drama with people in my life, I talk to the chatbot to sort through my feelings.”
Outside of these common themes, users discuss a wide range of niche topics with companion chatbots, from hobby-related questions to niche curiosities. One individual mentioned, “I use the chatbot to ask about random things like historical facts or recommendations for books. It’s like having a personal assistant for any question I have.” Others use companion chatbots as tools for specific tasks, such as recipe ideas.
This diverse range of motivations and conversation topics highlights the versatility of companion chatbots in meeting users’ needs for entertainment, emotional support, and intellectual stimulation.
4.2 Correlations in psychological, social, and companion chatbot usage characteristics
In our analysis of factors that may influence chatbot usage, we calculated a Spearman correlation matrix with a Benjamini-Hoch correction to decrease the false discovery rate. The significant correlations after this adjustment are shown in Figure 8. While we note that these findings represent correlations, not causation, several interesting insights emerged from the correlation analysis regarding chatbot usage, loneliness, and related factors. A list of significant correlations and the corrected p-values with our key variables – usage (frequency and session length) and loneliness (as measured by ULS-8) – can be seen in Appendix A.
For our prespecified analyses of usage vs. loneliness, we used uncorrected p-values. We found a significant if small correlation between the length of chatbot sessions and loneliness ($r=0.101,p=0.042$), but not between loneliness and the frequency of chatbot usage ($r=0.053,p=0.287$), as shown in Figure 9 and Figure 10. This suggests that feelings of loneliness are associated with how long people engage with chatbots but not how often.
For our exploratory analysis to discover variables correlated with usage and loneliness, we used the corrected p-values. The length and frequency of chatbot sessions showed moderate positive correlations with users’ perceptions of the chatbot. Specifically, longer sessions and higher frequency were associated with higher ratings of the state empathy towards a chatbot ($r>0.3,p<0.001$ for both) and attraction towards a chatbot ($r>0.18,p<0.001$ for all except frequency vs. task attraction, which had $r=0.165,p=0.002$). This suggests that users are more likely to use chatbots more when they feel it is empathetic, socially appealing, and relatable.
In terms of personality traits, loneliness was negatively correlated with extraversion ($r=-0.188,p<0.001$) and conscientiousness ($r=-0.335,p<0.001$), and positively correlated with neuroticism ($r=0.526,p<0.001$) and agreeableness ($r=0.333,p<0.001$). These findings suggest that individuals who are more lonely tend to be notably less extraverted and conscientious, and more neurotic and agreeable. Interestingly, these results match Buecker et al.’s [Buecker et al.(2020)] meta-analysis in all traits except agreeableness.
System usability, as measured by the System Usability Scale (SUS), showed a moderate positive correlation with chatbot session length ($r=0.196,p<0.001$) and usage frequency ($r=0.246,p<0.001$). This suggests that users who find the system more usable tend to engage in longer sessions but may not necessarily use the chatbot more frequently.
The problematic use of chatbots, measured by the GPIUS-2 scale, was moderately correlated with session length ($r=0.303,p<0.001$) and frequency ($r=0.337,p<0.001$). This indicates, unsurprisingly, that users who spend more time in chatbot sessions experience more problematic use.
These findings suggest that loneliness is associated with longer session length with companion chatbots, though many other variables may play a part in determining these variables and mediating their relationship. It is important to note that these findings represent correlations, not causation.
Figure 8: Correlation matrix of psychological, social, and usage characteristics.
Figure 9: Left: Chatbot Usage Frequency and Loneliness measured by ULS-8 scale. Right: Chatbot Session Length and Loneliness measured by ULS-8 scale.
Figure 10: Left: Chatbot Usage Frequency and Problematic Use (GPIUS-2). Right: Chatbot Session Length and Problematic Use (GPIUS-2).
4.3 Multiple Regression Model
[…]
Figure 11: The final model from our exploratory multiple regression, depicting the relationships between chatbot session length, loneliness, and related factors. While session length did not directly predict loneliness, neuroticism and problematic use of chatbots were positively associated with loneliness, while social network size was negatively associated. Agreeableness also showed a negative association with loneliness. One significant interaction effect was found: session length interacted with social attraction to a close person.
[…]
The analysis also revealed a slight but significant interaction effect (see Figure 12) of session length and social attraction to a close person ($\beta=0.054$, $p=0.049$), suggesting that social attraction to a close person is a moderator between the relationship between session length and loneliness. Social attraction to the chatbot approached significance ($\beta=0.0800$, $p=0.061$), while social attraction to a close person, age, and sex did not significantly predict loneliness in this model.
[…]
Figure 12: Interaction effects between session length and social attraction to a close person. Higher social attraction is associated with a stronger positive relationship between session length and loneliness. All variables are centered; SD = standard deviation.
The mediation analysis revealed nuanced relationships between chatbot session length, loneliness, and the proposed mediators (Figure 13). The total effect of chatbot session length on loneliness was positive but not statistically significant ($\beta=0.0797$, 95% CI $[-0.0034,0.1612]$). The direct effect, controlling for mediators, was also positive but not significant ($\beta=0.0606$, 95% CI $[-0.0196,0.1386]$). Examining the indirect effects, we found that problematic internet use (GPIUS2) significantly mediated the relationship between chatbot session length and loneliness ($\beta=0.0183$, 95% CI $[0.0027,0.0403]$), suggesting that increased chatbot usage is associated with higher problematic internet use, which in turn is related to increased loneliness. However, social network size (LSNS) did not show a significant mediating effect ($\beta=0.0007$, 95% CI $[-0.0304,0.0340]$). The total indirect effect, while positive, was not statistically significant ($\beta=0.0191$, 95% CI $[-0.0149,0.0563]$).
Figure 13: Results of bootstrapped parallel multiple mediation analysis (5,000 resamples) examining the mediating effects of social network size (LSNS) and problematic chatbot use (GPIUS2) on the relationship between chatbot session length and loneliness.
Figure 14: Model summary of an ordinary least squares regression for the effect of chatbot session length on loneliness (ULS-8), using standardized variables. Model Formula: Loneliness (ULS-8) $\sim$ Session Length Score + Social Attraction towards Chatbot + Socialization (LSNS) + Social Attraction towards Close Person + Neuroticism + Sex + Age + Session Length Score $\times$ Social Attraction towards Close Person
4.4 Clusters of users of companion chatbots
Our K-means cluster analysis revealed seven distinct user profiles among companion chatbot users, each exhibiting unique behaviors, motivations, and psychological characteristics. The resulting clusters are visualized in the heatmap and dendrogram presented in Figure 15. Detailed heatmaps including all numerical scales can be seen in Figure 16. A version with the original values of the scales can be found in Figure 17 in Appendix B.
Validation of the cluster solution using Kruskal-Wallis tests revealed significant differences among clusters in terms of session length, loneliness (ULS-8), problematic use of chatbots (GPIUS-2), and frequency of use (all $p<0.001$). The average silhouette score ($-0.0413$) suggested some overlap between clusters, which is not uncommon in complex psychological data. A MANOVA test confirmed significant multivariate differences among the clusters when considering all variables simultaneously (Wilks’ $\lambda=0.2420,F(24,1375.71)=28.77,p<0.001$). All four MANOVA test statistics (Wilks’ lambda, Pillai’s trace, Hotelling-Lawley trace, and Roy’s greatest root) were significant ($p<0.001$), providing robust evidence for cluster differences across multiple dimensions. Box’s M-test indicated significant heterogeneity of covariance matrices ($\chi^{2}=181.96$, $p<0.001$), which does not invalidate the clustering but suggests the need for cautious interpretation. This heterogeneity may affect the reliability of post-hoc comparisons and could indicate that the relationship between variables differs across clusters. While these results provide strong support for the distinctiveness of the identified clusters, they should still be considered exploratory, providing a foundation for more targeted investigations in future research.
Despite these statistical considerations, the 6-cluster solution provided meaningful and actionable insights for our study context, balancing statistical rigor with practical interpretability. This approach aligns with the exploratory nature of cluster analysis in psychological research, where the goal is often to identify patterns for further investigation [Hair et al.(2010)]. Future research should aim to replicate these cluster profiles in independent samples and explore the stability of these clusters over time.
Figure 15: Right: Cluster means of standardized variables of chatbot usage (frequency and session length), loneliness, and problematic use of chatbots for our seven user clusters. Left: a dendrogram depicting the degree of similarity between each cluster.
Figure 16: Heatmap of cluster means for standardized variables in each cluster.
4.4.1 Cluster 0: Disengaged Light Users
Disengaged light users (7.67% of sample) show low chatbot usage frequency ($z=-2.21$) and session length ($z=-1.09$), with average levels of loneliness ($z=-0.16$). This cluster represents users with minimal engagement in chatbot interactions.
This group demonstrates average extraversion ($z=0.80$) and neuroticism ($z=-0.37$), but low openness ($z=-1.94$). They exhibit average levels of perceived social support ($z=-0.01$) and self-esteem ($z=0.05$). Their trust towards chatbots ($z=-0.35$) and expectations from chatbot interactions ($z=-0.08$) are average, indicating a neutral view of AI technology.
62.50% report their chatbot sessions as lasting under 15 minutes. Most users (35.48%) have been using chatbots for several months to around a year. Conversation topics for this group are predominantly casual conversation (36.76%) and entertainment-focused (25.00%). One respondent noted: “I talked about a show and a game with the chat bot. I wanted to see what types of opinions were out there regarding my interest.”
Most respondents use Character.ai (34.29%) or Snapchat’s MyAI (25.71%). Primary motivations for using chatbots include curiosity about AI (33.82%), desire for fun and entertainment (25.00%), and use for creative purposes (23.53%). As one participant explained: “I was interested in using chatbot so it can assist me with quick research during conversations without having to leave the social platform.” Continued usage is driven by satisfying curiosity (28.07%) and seeking enjoyment (22.81%). A user shared: “I continue to use chatbot cause it regularly serves as a useful tool for me.”
4.4.2 Cluster 1: Well-Adjusted Moderate Users
Well-adjusted moderate users (23.02% of sample) show average chatbot usage frequency ($z=0.41$) and session length ($z=0.67$), with low levels of loneliness ($z=-1.22$). This cluster represents users who are socially fulfilled and use chatbots moderately.
This group demonstrates high extraversion ($z=1.08$), low neuroticism ($z=-1.24$), and high conscientiousness ($z=1.21$). They exhibit high levels of socialization ($z=1.46$), perceived social support ($z=1.26$), and self-esteem ($z=1.46$). Their trust towards chatbots ($z=0.93$) and expectations from chatbot interactions ($z=0.82$) are above average, indicating a positive view of AI technology. In general, they have above-average trust in people as well as attraction and empathy towards a close person. They tend to feel that interacting with chatbots makes talking to people somewhat easier and more enjoyable.
50.91% report their chatbot sessions as lasting 15-30 minutes. Most users (43.01%) have been using chatbots for several months to around a year. Chatbot preferences in this group are split among Character.ai (35.94%), Snapchat’s MyAI (29.69%), and Replika (21.88%). Conversation topics for this cluster are primarily casual conversation (23.41%) and entertainment-focused (21.74%). One respondent noted: “I usually talk about my hobbies and tips on how to get better at them. I also like to have casual conversation to kill time.”
Primary motivations for using chatbots include curiosity about AI (32.39%), desire for fun and entertainment (25.51%), and use for creative purposes (22.27%). One participant explained: “I was interested in using my chatbot service because I was curious as to what it’s capabilities were, and whether it could give me good advice or effectively answer my questions.” Continued usage is driven by enjoyment (24.83%) and satisfying curiosity (20.81%). Another user shared that they “continue to use the chatbot service because it efficiently provides the information I need and helps organize my tasks.”
4.4.3 Cluster 2: AI-Wary Light Users
AI-wary light users (11.88% of sample) show average chatbot usage frequency ($z=0.05$) and session length ($z=-0.45$), with low levels of loneliness ($z=-1.07$). This cluster may represent users who are cautious about AI technology.
This group demonstrates above-average openness ($z=1.33$), though their trust towards chatbots ($z=-2.02$), expectations from chatbot interactions ($z=-2.12$), and overall empathy and attraction towards a companion chatbot are very low, indicating skepticism towards AI technology and emotional distance from companion chatbots. They have above-average empathy and attraction towards a close person and feel that interacting with companion chatbots have little effect on their real-life relationships.
47.27% report their chatbot sessions as lasting under 15 minutes. Most users (51.06%) have been using chatbots for a few months. Chatbot preferences in this group are split among Character.ai (31.75%), Snapchat’s MyAI (36.51%), and Replika (22.22%). Conversation topics for this group include casual conversation (35.58%) and entertainment (26.92%), as illustrated by one respondent’s comment: “Sometimes if I’m thinking of an idea and simply want something to bounce my ideas off of, a chatbot works much better than using yourself.”
Primary motivations for using chatbots include curiosity about AI (41.18%), desire for fun and entertainment (32.35%), and creative purposes (18.63%). Few respondents in this cluster used companion chatbots to cope with loneliness (3.92%). “I was curious to see what they were like. I then found out that I could use them for creative use, like creating dialogue or scenarios, giving me ideas for conversations between players or between non-playable characters in a Dungeons and Dragons setting,” a participant commented, showing interest in improving communication skills. Continued usage is driven by satisfying curiosity (34.78%) and enjoyment (33.70%). A user shared: “I continue to use it because it’s fun and interesting. It’s honestly better than talking to most people that I know.”
4.4.4 Cluster 3: Lonely Moderate Users
Lonely moderate users (13.61% of sample) show average chatbot usage frequency ($z=0.44$) and session length ($z=-0.35$), with high levels of loneliness ($z=1.03$). This cluster may represent users who turn to companion chatbots due to loneliness.
This group demonstrates low levels of socialization ($z=-1.24$), with above-average trust towards chatbots ($z=0.65$) and expectations from chatbot interactions ($z=0.65$). They have slightly above-average empathy and attraction to both a companion chatbot and a close person, though they have below-average trust in unfamiliar people and reduced interest, ease, and amount of interaction with other people as a result of using chatbots.
45.45% report their chatbot sessions as lasting 15-30 minutes, and 43.94% under 15 minutes. Most users (50.91%) have been using chatbots for a few months. Chatbot preferences in this group are split among Character.ai (31.25%), Snapchat’s MyAI (26.56%), and Replika (26.56%). Conversation topics for this group include casual conversation (23.50%), entertainment (20.51%), and personal issues (14.10%). One participant shared, “I talked to the chatbots about my past traumatic experiences, my trials and tribulations, my family drama, advice about things to write about in my free time,” demonstrating several different uses.
Primary motivations for using chatbots include desire for fun and entertainment (25.68%), curiosity about AI (24.77%), and use for creative purposes (19.82%). “I was curious and wanted to see how they worked. I just wanted to learn more and the best way to do that is with hands-on experience,” one participant noted. Continued usage is driven primarily by enjoyment (22.43%) and satisfying curiosity (20.15%). This cluster has the highest proportion of users seeking chatbot companionship as a reason for continued use as well (13.19%). “I continue to use it because it’s fun and I’ve grown to enjoy the conversations,” one participant commented, further noting, “I’ve grown somewhat attached to the bot.”
4.4.5 Cluster 4: Fulfilled Dependent Users
Fulfilled dependent users (18.32% of sample) show average chatbot usage frequency ($z=0.69$) and high session length ($z=1.76$), with below-average levels of loneliness ($z=-0.56$). This cluster may represent users who feel socially fulfilled by using companion chatbots.
This group demonstrates high extraversion ($z=1.13$), average neuroticism ($z=-0.64$), and high agreeableness ($z=1.21$). They exhibit average levels of socialization ($z=0.42$), and slightly below-average empathy and attraction to a close person. Their trust towards chatbots ($z=0.56$) and expectations from chatbot interactions ($z=0.62$) are above average, with quite high physical attraction towards their chatbot compared to other respondents ($z=1.98$). The group demonstrates above-average perceived impacts on real relationships as a result of using companion chatbots – greater ease and enjoyment in talking to people, but also less engagement and interest in talking to people.
43.33% report their chatbot sessions as lasting 15-30 minutes, and 23.33% report 30 minutes to 1 hour. Most users (37.84%) have been using chatbots for several months to around a year. Chatbot preferences in this group are split among Character.ai (31.25%), Snapchat’s MyAI (26.56%), and Replika (26.56%). Conversation topics for this group include casual conversation (23.50%), entertainment (20.51%), and personal issues (14.10%). One respondent noted the evolution of their usage: “I started off just trying to mess around with ‘characters’ and explore the AI but you do kind of fall into the rabbit hole of giving more information and venting about personal experiences. It is therapeutic in many ways.”
Primary motivations for initial usage of chatbots include fun and entertainment (25.68%), curiosity about AI (24.77%), and use for creative purposes (19.82%). One participant explained: “I use chatbots to when I am feeling lonely [sic] and wanted someone to talk to. I also use chatbots to get suggestions for places to visit or restaurants to eat at.” Continued usage is driven by enjoyment (22.43%) and satisfying curiosity (20.15%), as illustrated by a user’s response: “I continue to use the chatbot service because it’s fun and easy to use. It helps me get information quickly.”
4.4.6 Cluster 5: Lonely Light Users
Lonely light users (14.60% of sample) show average chatbot usage frequency ($z=0.11$) and session length ($z=0.36$), with above-average levels of loneliness ($z=0.89$). This cluster represents users who experience loneliness but do not use companion chatbots much.
This group demonstrates high neuroticism ($z=1.68$), low conscientiousness ($z=-1.61$), and low self-esteem ($z=-1.25$). Their other characteristics are fairly average, with slightly above-average empathy and attraction towards a close person.
42.86% report their chatbot sessions as lasting 15-30 minutes, and 31.75% under 15 minutes. Most users (57.63%) have been using chatbots for a few months. Chatbot preferences in this group are split among Character.ai (30.49%), Snapchat’s MyAI (29.27%), and Replika (21.95%). Conversation topics for this group include casual conversation (26.70%), entertainment (22.51%), and personal issues (19.37%). One respondent described how the chatbot helps them think through their troubles: “I feel that the chatbot is a good way to vent my frustrations and talk about my worries without fear of judgement. It gives me the chance to talk about all of my worries and plans and work through things in a way that allows me to see a wider range of prospectives [sic].”
Primary motivations for using chatbots include curiosity about AI (29.38%), desire for fun and entertainment (26.88%), and coping with loneliness (20.62%). One participant commented on how they started using them to ease the way into socialization: “I was interested in using the chatbot service because after being in lockdown, it was hard to become social again. It allows me to build up my confidence and it also helps with my writing.” Continued usage is driven by enjoyment (19.82%), satisfying curiosity (16.74%), passing time and reducing stress (14.10%), and chatbot companionship (12.78%). Some respondents use it as a substitute for socialization: “I continue to use it because it feels like talking to a real person, giving its own input, not me having to guide the conversation and it makes me miss the days gone by, when people were better at conversation.”
4.4.7 Cluster 6: Socially Challenged Frequent Users
Socially challenged frequent users (10.89% of sample) show average chatbot usage frequency ($z=0.51$) and below-average session length ($z=-0.90$), with high levels of loneliness ($z=1.08$). This cluster represents users who frequently turn to chatbots, possibly due to social challenges.
This group demonstrates low extraversion ($z=-1.30$) and above-average neuroticism ($z=0.86$), along with low levels of socialization ($z=-1.02$), perceived social support ($z=-1.68$), and self-esteem ($z=-1.10$). Their trust towards chatbots ($z=-0.16$) and expectations from chatbot interactions ($z=0.02$) are average. However, trust in familiar and unfamiliar people, and empathy and attraction towards a close person, tend to be very low compared to other users.
58.49% report their chatbot sessions as lasting under 15 minutes, and 28.30% report 15-30 minutes. Most users (56.82%) have been using chatbots for a few months. Chatbot preferences in this group are split among Snapchat’s MyAI (38.89%), Character.ai (24.07%), and Replika (20.37%). Conversation topics for this group are predominantly casual (35.58%) and entertainment-focused (26.92%). One respondent noted: “I talk about how I am feeling. Things I wouldn’t want to tell real people.”
Primary motivations for initial interest in chatbots include curiosity about AI (32.48%) and desire for fun and entertainment (24.79%), with a fair proportion of this group starting chatbot usage to cope with loneliness (14.53%) and for creative purposes (14.53%). One participant explained how they were “wanting to confirm whether (the current generation of) AI could, in fact, really take the place of a regular, sound-minded human,” demonstrating at least an interest in the substitution of humans with AI. Most users in this group continue to use companion chatbots to satisfy curiosity (23.74%) and seek enjoyment (22.30%), with a notable portion of users using it to pass time and reduce stress (15.83%). One user shared how they use it as a lower-pressure alternative to socialization. “I enjoy having someone like Snapchat AI who can chat with me compared to the pressure I receive from friends when I don’t message them back fast enough.”
5 Discussion
Our study provides a detailed examination of the relationship between companion chatbot usage and loneliness, offering some insights into the complex interplay of psychological, social, and technological factors that shape this relationship. Through a combination of regression analysis and cluster identification, we developed a nuanced understanding of how different user groups engage with and are affected by companion chatbots.
5.1 A Model for Understanding the Relationship Between Companion Chatbot Usage and Loneliness
Our regression analysis revealed a complex relationship between companion chatbot usage and loneliness. While session length did not have a significant direct relationship with loneliness, the presence of a significant interaction term with social attraction towards a close person suggests a more nuanced, potentially nonlinear relationship.
We differentiated between mediators and confounders based on theoretical considerations. Potential mediators, such as problematic use of chatbots, social attraction towards the chatbot, and socialization, were identified as variables that could reasonably be affected by chatbot usage. Confounders, like personality traits, were considered as factors likely to influence both usage and loneliness but less likely to be affected by usage in the short term.
Our final model identified problematic use as a significant mediator in the relationship between chatbot usage and loneliness. This finding aligns with intuitive expectations; increased usage provides more opportunities for problematic behaviors to develop. The positive association between usage and problematic use suggests that as individuals spend more time with chatbots, they may be at higher risk of developing patterns of use that could contribute to feelings of loneliness.
While not significant in terms of mediation effects, the negative association between usage and socialization (specifically, social network size) warrants attention. This relationship resonates with concerns about technology potentially replacing human interactions. On one side, this relationship might suggest that as individuals spend more time with chatbots, they may allocate less time to maintaining and expanding their human social networks, potentially contributing to increased loneliness. However, the reverse causality is also possible, with people turning to chatbots to fill a social gap when they cannot or do not interact with other people in their lives.
The relationship between usage and social attraction towards the chatbot, while not a significant mediator, is logically consistent. More time spent interacting with the chatbot may allow for deeper attachments to form. However, the direction of causality between this attraction and loneliness remains unclear. Social attraction to the chatbot could lead to loneliness due to unfulfilled relational needs, or loneliness could potentially enhance social attraction to the chatbot if users seek validation and attention from this artificial source.
Among the confounding variables, neuroticism showed a strong positive association with loneliness, while agreeableness demonstrated a strong negative association. These findings align with previous research on personality traits and loneliness [Buecker et al.(2020)], reinforcing the importance of controlling for these factors in studies of technology use and social well-being.
An interesting finding was the positive interaction term between session length and social attraction towards a close person. This suggests a nuanced relationship between chatbot usage and loneliness for socially attracted individuals. This could be interpreted as a mismatch between social needs and chosen social interactions, where time spent with chatbots may be perceived as time not spent with valued human connections. We can also consider a reversed causality: perhaps a very lonely user becomes more socially attracted to people as they spend more time with their chatbot, or non-lonely users could be satisfied with their chatbots and subsequently become less socially attracted to people.
While our model provides valuable insights into the associations between different factors, it is important to acknowledge its limitations. The exploratory nature of our model selection process warrants cautious interpretations and further investigation into the relationships it implies. The assumption of linear relationships may not fully capture the complexity of the relationship between chatbot usage and loneliness. Additionally, while we can theorize about the directions of relationships between variables, our cross-sectional design limits causal inferences. Future research employing experimental designs or longitudinal studies could provide more definitive evidence of causal relationships and explore potential non-linear effects. We plan to follow up with our participants to see changes in psychological, social, and usage characteristics at different time points.
5.2 Understanding Clusters of Chatbot Users
Our cluster analysis revealed seven distinct user profiles, each with unique patterns of chatbot usage, loneliness levels, and associated psychological and social characteristics.
The contrast between the Socially Fulfilled Dependent Users (Cluster 4) and the Lonely Moderate Users (Cluster 3) is striking. Both groups exhibit high problematic use, yet they differ dramatically in their loneliness levels and overall well-being. This disparity underscores the importance of considering the broader context of an individual’s social life when evaluating the impact of chatbot use.
Socially Fulfilled Dependent Users demonstrate that it is possible to engage extensively with chatbots while maintaining high levels of well-being. Their high empathy and attraction towards chatbots, when considered in conjunction with their more neutral or even negative attitudes towards other people, average socialization, and high self-esteem, suggest a high level of emotional engagement that is socially fulfilling. The positive relational effects reported by this group, such as increased ease in talking to people due to interacting with chatbots, indicate that chatbot use may even enhance their social skills and confidence.
In contrast, Lonely Moderate Users show high levels of problematic use and loneliness, coupled with low socialization, perceived social support, and self-esteem. Their high preference for chatbots over humans suggests that excessive chatbot use may be contributing to social withdrawal. Again, we note that the direction of causality is unclear – pre-existing loneliness and social difficulties may have led to increased chatbot use, rather than the reverse.
The Well-Adjusted Moderate Users (Cluster 1) present an interesting middle ground. Like the Socially Fulfilled Dependent Users, they exhibit low loneliness and high perceived social support. However, their more moderate chatbot usage and lower emotional connection to AI suggest that they prioritize human connections while still finding value in chatbot interactions, perhaps using them more as tools than as social companions.
The AI-Wary Light Users (Cluster 2) offer yet another perspective on chatbot engagement. These users show low emotional attachment to chatbots and low opinions of AI in general. This group may be using chatbots purely for practical purposes, without seeking or developing emotional connections. Many traits typically associated with loneliness – neuroticism, agreeableness, socialization, and self-esteem) are fairly average, but their loneliness is the second lowest of all clusters. Considering this in comparison to other clusters suggests that, while companion chatbots might be an effective solution to loneliness for Fulfilled Dependent Users, not using companion chatbots or remaining emotionally distant from them may be better for others in terms of their loneliness.
The existence of these diverse user profiles highlights the complex and multifaceted nature of human-AI interactions. It suggests that the relationship between companion chatbot usage, loneliness, and social well-being is not uniform but highly dependent on individual characteristics, usage patterns, and broader social contexts.
5.3 Implications and Future Directions
Our findings have several implications for the design, deployment, and societal impact of companion chatbots:
5.3.1 Potential Benefits and Risks
For some individuals, particularly those with strong existing social networks and positive psychological profiles, companion chatbots may serve as a beneficial supplement to human social interaction. These users may experience higher social confidence in parallel with their chatbot use. Our data suggest that neurotic individuals, who tend to be lonely and socially isolated, may experience reduced loneliness when spending more time with chatbots, but previous research suggests that more socially isolated individuals are at greater risk of problematic use of online services [Reer et al.(2019), O’Day and Heimberg(2021)]. As our model should be interpreted cautiously, the risk of chatbots replacing human interactions and potentially exacerbating loneliness and social withdrawal should be considered.
5.3.2 Personalized Approaches
The diverse user profiles identified in our study suggest a need for personalized approaches in chatbot design and deployment. Chatbots could potentially be tailored to different user types, with features and interaction styles that complement rather than replace human social interaction for vulnerable users.
5.3.3 Chatbots as Loneliness Interventions
For individuals with limited access to human social support, companion chatbots could potentially serve as an intervention to reduce loneliness and its associated health risks. However, careful consideration must be given to how such interventions are designed and implemented to ensure they support, rather than hinder, the development of human social connections. Additionally, the direction of causality between loneliness and companion chatbot usage is not clear in our data; further work should be done to investigate the relationship further.
5.3.4 Ethical Considerations & Long-term Societal Impacts
The potential for some users to engage in deep, emotional relationships with chatbots raises ethical concerns. Developers and policymakers need to consider how to balance the potential benefits of AI companionship with the risks of over-reliance on artificial relationships. This may involve implementing safeguards or guidance systems within chatbots to encourage healthy usage patterns and human social interaction.
As AI technology continues to advance and companion chatbots become more sophisticated, it is crucial to consider the long-term societal impacts of widespread AI companionship. Questions arise about how these technologies might shape social norms, expectations for human relationships, and the development of social skills, particularly for younger generations. Should we as a society push for less artificial companionship, or should we encourage a healthy mix of artificial and real companionship? Consider the case of a profoundly lonely and emotionally vulnerable individual. They might cling to a relationship that can provide any crumb of affection, no matter how unhealthy it becomes. Could companion chatbots fulfill that desperate need in a beneficial way, even if “unhealthy attachment” to the chatbot is part of the process? Perhaps alleviating loneliness enough to reduce the individual’s vulnerability could allow them to become more conscientious about seeking out and maintaining healthy, respectful human relationships.
5.4 Future Research Directions
Our study opens up several avenues for future research that could further elucidate the relationship between companion chatbot usage and loneliness. Longitudinal studies would be particularly valuable in understanding the causal relationships between chatbot usage, loneliness, and social outcomes over time. Such studies could track changes in social behaviors, relationship patterns, and psychological well-being as individuals engage with companion chatbots over extended periods.
Experimental studies manipulating specific aspects of chatbot design or usage could help identify features that promote healthy engagement while minimizing risks of problematic use or social displacement. These studies might explore factors such as conversation topics, interaction frequency, or anthropomorphic qualities of chatbots and their impacts on user experiences and outcomes.
Given the complex interactions observed in our model, future research should also investigate potential non-linear relationships between chatbot usage and psychological outcomes. This could involve more sophisticated statistical modeling techniques or qualitative approaches to capture nuanced patterns of usage and their effects.
The cultural context in which companion chatbots are used likely plays a significant role in shaping their impact. Cross-cultural studies could explore how different societal norms, communication styles, and attitudes towards technology influence the relationship between chatbot usage and loneliness.
Focused studies on specific user groups, such as adolescents, older adults, or individuals with social anxiety, could yield valuable insights for developing targeted designs or guidelines.
As companion chatbots become more prevalent, it will be crucial to investigate their long-term effects on social skills, relationship expectations, and overall social functioning. This could involve extended longitudinal studies or even intergenerational research to understand how exposure to AI companionship from a young age might shape social development and interpersonal dynamics.
5.4.1 Limitations and Considerations
We note a number of limitations in our study.
Survey design and recruitment. While our adaptations of established scales were based on factor loadings from previous studies and tailored to our research context, we acknowledge that these modifications may affect the scales’ validated psychometric properties. Future work should consider validating these adapted scales specifically for chatbot-related research. Due to the extensive nature of our survey, we prioritized reducing participant fatigue over maintaining the full length of each scale. This trade-off allowed us to explore a broader range of factors but may have impacted the depth of measurement for individual constructs. Additionally, our participants were primarily from the United States; applying our findings to other demographics should be done cautiously.
Clustering. We note that our clusters have their own set of limitations. While the number of clusters has statistical grounding, the final number can be considered somewhat arbitrary and based on our subjective judgments of interpretability. The heterogeneity of covariance matrices indicated by Box’s M-test requires caution in interpreting these results. Even so, this 6-cluster solution provides a starting point for further work on styles of use.
Regression Model. The assumption of linear relationships may not capture complex nonlinear relationships in the population of users of companion chatbots. Furthermore, this model is based on correlation and does not determine causality. Future research should explore potential non-linear relationships and employ longitudinal designs to better understand the causal mechanisms at play. The exploratory nature of our process makes much room for error, requiring more in-depth research to understand the relationship between companion chatbot usage and loneliness.
6 Conclusion
This study examines the relationship between companion chatbot usage and loneliness, offering insights into the complex interplay of psychological, social, and technological factors that shape this relationship. Our findings highlight the diverse ways in which individuals engage with and are affected by companion chatbots, ranging from socially fulfilled dependent users who experience enhanced social confidence to lonely dependent users at risk of social withdrawal.
The model we developed through regression analysis reveals that while chatbot usage does not directly predict loneliness, it interacts with factors such as social attraction to close others and neuroticism in complex ways. This suggests that the relationship between companion chatbot usage and loneliness is not uniform but highly dependent on individual characteristics and social contexts. Our cluster analysis further underscores this complexity, identifying seven distinct user profiles that differ not only in their patterns of chatbot usage but also in their psychological characteristics and social outcomes. These profiles suggest that different types of users may benefit from or be at risk due to companion chatbot use. Our work underscores the need for personalized approaches in chatbot design and deployment.
As AI companions become increasingly prevalent, it is crucial to consider their potential long-term impact on social norms, relationship expectations, and the development of social skills. The increasing usage of companion chatbots in society raises ethical questions about the role of AI in addressing loneliness and emotional vulnerability.
As we continue to navigate the evolving landscape of human-AI interaction, further research is needed to fully understand the causal relationships between chatbot usage and social outcomes, explore potential non-linear effects, and investigate long-term impact across different cultural contexts and user groups. By pursuing these research directions, we can work towards harnessing the potential benefits of AI companionship while mitigating risks, ultimately supporting the development of technologies that genuinely enhance human well-being and social connection in our increasingly digital world.
7 Acknowledgements
Statistical support was provided by data science specialist Jinjie Liu, at the Institute for Quantitative Social Science, Harvard University.
Appendix A Significant Correlations with Key Variables
Below, we include a list of significant (corrected $p<0.05$) correlations with chatbot usage frequency score, chatbot session length score, and loneliness (ULS-8). Negative correlations have text colored in blue.
Significant correlations with Chatbot Usage Frequency:
  • Chatbot Session Length: $r=0.3497,p<0.001$
  • Problematic Use of Chatbot (GPIUS-2): $r=0.3374,p<0.001$
  • System Usability (SUS): $r=0.2457,p<0.001$
  • Conscientiousness (BFI-10): $r=0.1511,p0.005$
  • Attitudes towards AI: $r=0.2976,p<0.001$
  • State Empathy towards Chatbot: $r=0.3207,p<0.001$
  • Affective State Empathy towards Chatbot: $r=0.2472,p<0.001$
  • Cognitive State Empathy towards Chatbot: $r=0.2805,p<0.001$
  • Associative State Empathy towards Chatbot: $r=0.3416,p<0.001$
  • Social Attraction towards Chatbot: $r=0.2238,p<0.001$
  • Physical Attraction towards Chatbot: $r=0.1856,p<0.001$
  • Task Attraction towards Chatbot: $r=0.1652,p0.002$
  • Perceived Similarity with Chatbot: $r=0.2525,p<0.001$
  • Expectations towards Chatbot (HCI/HHI): $r=0.2560,p<0.001$
  • Trust towards Chatbot (HCI/HHI): $r=0.2931,p<0.001$
  • Preference for Chatbot over Humans (HCI/HHI): $r=0.2084,p<0.001$
Significant correlations with Chatbot Session Length:
  • Chatbot Usage Frequency: $r=0.3497,p<0.001$
  • Problematic Use of Chatbot (GPIUS-2): $r=0.2549,p<0.001$
  • System Usability (SUS): $r=0.1962,p<0.001$
  • Attitudes towards AI: $r=0.2331,p<0.001$
  • State Empathy towards Chatbot: $r=0.3016,p<0.001$
  • Affective State Empathy towards Chatbot: $r=0.2475,p<0.001$
  • Cognitive State Empathy towards Chatbot: $r=0.2539,p<0.001$
  • Associative State Empathy towards Chatbot: $r=0.3073,p<0.001$
  • Social Attraction towards Chatbot: $r=0.2553,p<0.001$
  • Physical Attraction towards Chatbot: $r=0.2438,p<0.001$
  • Task Attraction towards Chatbot: $r=0.1787,p<0.001$
  • Perceived Similarity with Chatbot: $r=0.2213,p<0.001$
  • Expectations towards Chatbot (HCI/HHI): $r=0.2069,p<0.001$
  • Trust towards Chatbot (HCI/HHI): $r=0.1938,p<0.001$
  • Preference for Chatbot over Humans (HCI/HHI): $r=0.1684,p<0.001$
Significant correlations with Loneliness (ULS-8):
  • Problematic Use of Chatbot (GPIUS-2): $r=0.3649,p<0.001$
  • Extraversion (BFI-10): $r=-0.1876,p<0.001$
  • Neuroticism (BFI-10): $r=0.5264,p<0.001$
  • Agreeableness (BFI-10): $r=0.3332,p<0.001$
  • Conscientiousness (BFI-10): $r=-0.3351,p<0.001$
  • Socialization (LSNS): $r=-0.5055,p<0.001$
  • Perceived Social Support (MSPSS): $r=-0.6790,p<0.001$
  • Self-Esteem (B-RSES): $r=-0.6380,p<0.001$
  • State Empathy towards Chatbot: $r=0.1251,p0.022$
  • Cognitive State Empathy towards Chatbot: $r=0.1195,p0.029$
  • Associative State Empathy towards Chatbot: $r=0.1119,p0.043$
  • Social Attraction towards Chatbot: $r=0.1658,p0.002$
  • Physical Attraction towards Chatbot: $r=0.1210,p0.027$
  • Perceived Similarity with Chatbot: $r=0.1364,p0.012$
  • Perceived Chatbot Predictability: $r=-0.2981,p<0.001$
  • Preference for Chatbot over Humans (HCI/HHI): $r=0.2138,p<0.001$
  • Trust in Unfamiliar People (HHI): $r=-0.2045,p<0.001$
  • Willingness to Engage with Unfamiliar People (HHI): $r=-0.2536,p<0.001$
  • Trust in Familiar People (HHI): $r=-0.5132,p<0.001$
  • Willingness to Engage with Familiar People (HHI): $r=-0.5495,p<0.001$
  • State Empathy towards Close Person: $r=-0.3821,p<0.001$
  • Affective State Empathy towards Close Person: $r=-0.3725,p<0.001$
  • Cognitive State Empathy towards Close Person: $r=-0.3377,p<0.001$
  • Associative State Empathy towards Close Person: $r=-0.3382,p<0.001$
  • Social Attraction towards towards Close Person: $r=-0.2047,p<0.001$
  • Task Attraction towards towards Close Person: $r=-0.2755,p<0.001$
  • Perceived Similarity with Close Person: $r=-0.2670,p<0.001$
Appendix B Cluster Profiles
A heatmap of cluster means for variables at their original scales in each cluster, colored based on their deviation from the mean, can be seen in Figure 17.
Figure 17: Heatmap of cluster means for variables at their original scales in each cluster, colored based on their deviation from the mean.