### v2

# Chatbot Companionship: A Mixed-Methods Study of Companion Chatbot Usage Patterns and Their Relationship to Loneliness in Active Users

Abstract
Companion chatbots offer a potential solution to the growing epidemic of loneliness, but their impact on users’ psychosocial well-being remains poorly understood. This study presents a large-scale survey (n = 404) of regular users of companion chatbots, investigating the relationship between chatbot usage and loneliness. We develop a model explaining approximately 50% of variance in loneliness; while usage does not directly predict loneliness, we identify factors including neuroticism, social network size, and problematic use. We identify seven distinct clusters of users, from socially fulfilled dependent users to lonely moderate users. Different usage patterns can lead to markedly different outcomes, with some users experiencing enhanced social confidence while others risk further isolation. Our work contributes to the ongoing dialogue about the role of AI in social and emotional support, offering insights for developing more targeted and ethical approaches to AI companionship that complement rather than replace human connections.
Keywords Companion chatbots, loneliness, artificial intelligence, artificial agents, human-AI interaction, social support
Figure 1: Overview of the study’s methodology and research questions. A survey of 404 participants explored chatbot usage and its psychological impacts. RQ1 investigates the motivations and uses of chatbots through descriptive statistics and free responses. RQ2 examines the relationship between chatbot usage and loneliness using correlation and multiple regression analyses, identifying potential mediating and moderating factors. RQ3 focuses on identifying distinct user profiles through K-means cluster analysis, considering psychological and social factors. The results highlight the complexity of human-AI interactions and their varied impacts on emotional well-being.
1 Introduction
In an increasingly digital world, we are paradoxically more connected and isolated than ever. Despite unprecedented technological means to connect, loneliness remains pervasive, affecting approximately a third of individuals in industrialized countries [Loveys et al.(2019)]. This phenomenon has led to what the U.S. Surgeon General declared an “epidemic of loneliness”[Office of the Assistant Secretary for Health (OASH)(2023)]. The gravity of this issue is underscored by research indicating that chronic loneliness can increase mortality risk by 26-29%, comparable to smoking 15 cigarettes a day [Chawla et al.(2021)].
Artificial intelligence (AI) has emerged as a potential tool for addressing loneliness. Companion chatbots, which engage in complex conversations and maintain persistent relationships with users, have gained significant popularity, with platforms like Replika and Character.ai reporting over 30 million and 20 million users, respectively [Patel(2024), Shewale(2024)]. These AI companions offer users the opportunity for constant, judgment-free interaction. However, the impact of these artificial relationships on users’ psychological well-being and social behaviors needs further study.
Some research suggests that AI companions, even when used over the short term, decrease feelings of loneliness [Loveys et al.(2019), Gasteiger et al.(2021), De Freitas et al.(2024)]. However, concerns have been raised about the authenticity of these relationships, suggesting that they lower the bar for what people think relationships can be [Turkle(2011)]. This could lead to unrealistic expectations and further social withdrawal.
The nuanced nature of human-AI interaction in this context presents a critical gap in our understanding. Specifically, we lack comprehensive knowledge about factors influencing companion chatbots’ impact on loneliness and other psychological and social factors. As Xia et al. [Xia et al.(2024)] noted, artificial agents may complement human social interactions in some contexts while replacing them in others. This variability in outcomes underscores the need for a more comprehensive and granular analysis of user characteristics, usage patterns, and their relationships to psychological well-being.
To address this gap, we conducted a survey (n = 404) of regular companion chatbot users, with the following research questions:
  • RQ1: What motivates the use of companion chatbots, and what do people use them for?
  • RQ2: What is the relationship between companion chatbot usage and loneliness, and what factors mediate or moderate this relationship?
  • RQ3: Can we identify and characterize distinct user profiles based on usage patterns, loneliness levels, problematic use, and other psychological and social factors?
Our study (see Figure 1) aims to provide a more nuanced understanding of the impact of companion chatbots. By identifying the characteristics of users who experience psychosocial benefits from chatbot interactions (i.e. reduced loneliness and improved social skills) and those who may experience negative consequences (i.e. emotional dependence and social withdrawal), we hope to inform more targeted and ethical design practices. For example, more information could guide the development of customizable interfaces that emphasize features beneficial to specific user groups or implement safeguards for at-risk users. By examining how users’ mental models and expectations of AI companions influence outcomes, we can explore strategies for fostering healthy human-AI relationships that complement rather than replace human connections.
Through a combination of statistical analyses, including multiple regression and cluster analysis, along with qualitative assessment of user responses, we present the following contributions:
  • A model explaining approximately 50% of variance in loneliness, identifying problematic use as a key mediator and social attraction as a moderator in the relationship between chatbot usage and loneliness while accounting for personality traits and social network characteristics.
  • A typology of seven distinct companion chatbot user profiles demonstrating how similar usage patterns can lead to markedly different outcomes based on individual characteristics and social contexts.
  • An exploration of motivations and usage characteristics across different user types, revealing how technological exploration and entertainment dominate over social purposes despite marketing as social tools.
  • Design implications for creating more effective and ethically responsible AI companions, including the need for personalized approaches and built-in mechanisms to detect and intervene in potentially harmful usage patterns.
By addressing these questions, we aim to advance the field of human-computer interaction in the domain of AI companionship, providing insights that can guide the development of technologies that genuinely enhance human well-being and social connection in our increasingly digital world.
2 Background and Related Work
2.1 Loneliness and Social Support
Loneliness, defined as the subjective feeling of isolation regardless of one’s actual social network, has become a pressing global issue affecting approximately a third of individuals in industrialized countries [Loveys et al.(2019)]. Social isolation is the corresponding objective measure, often defined as a poorness in quantity and quality of contact with others. The U.S. Surgeon General’s declaration of loneliness as an epidemic in 2023 underscores the gravity of this issue [Office of the Assistant Secretary for Health (OASH)(2023)]. Chronic loneliness is associated with a significantly increased risk of mortality—approximately 26-29%, comparable to smoking 15 cigarettes a day or having an alcohol use disorder [Freedman and Nicolle(2020)].
Various risk factors contribute to social isolation and loneliness, including socio-demographic factors (e.g., greater age, low income), medical factors (e.g., cognitive impairment, sensory loss), and social factors (e.g., living alone, lack of transportation) [Freedman and Nicolle(2020)]. While there is evidence of loneliness peaking in adolescence in certain demographics [Yang and Victor(2011)], older populations are particularly vulnerable due to the accumulation of these risk factors and potentially lower digital literacy, which can limit their use of modern communication technologies [Gasteiger et al.(2021)].
Social support has been identified as a strong predictor of good physical and mental health, as well as resilience to stress. It can be defined as the support accessible to an individual via their social ties to others including individuals, groups, and a larger community [Ozbay et al.(2007)]. Importantly, the quality of relationships is a better predictor of health outcomes than the mere quantity of social connections [Wang et al.(2018), Schutter et al.(2021)].
2.2 Loneliness Coping Strategies and Interventions
Individuals employ various strategies to cope with loneliness, ranging from traditional approaches to digital solutions. Rokach (1998) identified six primary factors in coping with loneliness: Reflection and Acceptance (involving solitary reflection and acceptance of loneliness), Self-Development and Understanding (involving increased understanding of oneself through relationships and therapy), Social Support Network (increased social involvement), Distancing and Denial (involving unhealthy behaviors such as turning to drugs and self-harming), Religion and Faith (involving the feeling of belonging in religious communities), and Increased Activity (devoting more time to extracurricular activities) [Rokach and Brock(1998)]. A systematic review by Deckx et al. (2018) further categorized coping strategies into problem-focused and emotion-focused approaches, with problem-focused strategies consistently associated with lower levels of loneliness [Deckx et al.(2018)].
In recent years, digital platforms have emerged as commonly used coping mechanisms for loneliness, providing both social interaction and passive engagement. Individuals may use video platforms such as YouTube to cope with their loneliness – higher loneliness is associated with the formation of parasocial relationships, where individuals form one-sided relationships with content creators [Niu et al.(2021), de Bérail et al.(2019), Balcombe and Leo(2023)]. This is consistent with the parasocial compensation hypothesis, which postulates that lonely, socially isolated, and socially anxious individuals tend to be more engaged in parasocial relationships [Madison et al.(2016)].
Online communities and forums, such as Reddit, Facebook Groups, and Twitch, facilitate real-time communication and foster social connections across shared interests, allowing individuals to engage in virtual communities and potentially alleviate social isolation [Seering et al.(2019)]. Existing work suggests that seeking and offering social support are key activities in forums [Johnson et al.(2020)]. However, high levels of loneliness and social anxiety are associated with addiction and problematic use of these platforms, mediated by increased social comparison and fear of missing out [Reer et al.(2019), O’Day and Heimberg(2021), Burke et al.(2010)]. Interestingly, image-based media like Instagram may decrease loneliness due to the enhanced intimacy of image-sharing, while text-based media such as X (Twitter) does not show the same effect [Pittman and Reich(2016)].
There are, however, effective interventions for reducing loneliness. These typically fall into four categories, and can be delivered in-person or online: (1) modifying maladaptive social cognition (e.g., cognitive behavioral therapy, psychodynamic therapy, reminiscence, and mindfulness), (2) improving social support (e.g., involving care professionals in health and social care provision), (3) facilitating social interaction (e.g., clubs, shared interest groups, and videoconferencing), and (4) enhancing social skills (e.g., training programs and computer courses) [Rokach and Brock(1998), Gasteiger et al.(2021), Hickin et al.(2021)]. While these interventions can reduce loneliness, their efficacy varies across individuals. Hickin et al. [Hickin et al.(2021)] conducted a systematic review and meta-analysis of psychological interventions for loneliness, finding that such interventions are generally effective across the lifespan. However, they emphasize that the considerable heterogeneity in intervention effectiveness underscores the importance of personalized approaches. This finding aligns with the growing recognition that lonely individuals are a diverse group with varying needs [Cacioppo and Cacioppo(2018)], suggesting that interventions – including technological ones – must be tailored to individual circumstances for optimal impact.
2.3 Companion Chatbots
The emergence of companion chatbots, also known as virtual friends or digital companions, represents a novel approach to tackling loneliness [Wang et al.(2024), Goodings et al.(2024), Boine(2023), Chen et al.(2024), Shani et al.(2022), Brewer(2022)]. We refer to these as commercially available chatbots that employ generative AI to engage in complex conversations and maintain persistent relationships with users. These companion chatbots are accessible to a broad demographic, with several popular companion chatbot services (see Figure 2) having gained significant user bases:
Figure 2: Examples of companion chatbots: Replika, Character.ai, Snapchat’s My AI, Anima, and Pi.
  • Replika: An app featuring a generative AI chatbot designed to act as a friend, sibling, assistant, or romantic partner. Users interact with one chatbot that adaptively develops over time through continued conversations. Replika’s CEO reported over 30 million registered users worldwide in August 2024 [Patel(2024)], with its audience primarily aged 25-30 [Christopher(2023)].
  • Character.ai: Established in late 2022, this platform allows users to role-play using constructed personas and interact with chatbots based on fictional characters. With over 20 million users reported in March 2024, its audience primarily comprises young adults aged 18-24 [Shewale(2024)]. Notably, the sixth most popular character on the platform is “Psychologist,” with over 70 million messages shared in a year, indicating a significant trend of users seeking mental health support from AI [Tidy(2024)].
  • Snapchat’s My AI: A chatbot integrated into the Snapchat platform, powered by large language models like GPT-3.5 and GPT-4. It offers advice, answers questions, and assists with various tasks, demonstrating the integration of AI companions into mainstream social media platforms. Over 150 million individuals have used My AI since its launch [Hutchinson(2023)].
  • Anima: A chatbot service that enables role-playing conversations with different personas, allowing users to explore various scenarios and characters.
  • Pi: Developed by Inflection AI, Pi is an AI companion designed to engage in open-ended conversations on a wide range of topics. It aims to be a supportive, empathetic presence for users, offering advice, emotional support, and intellectual discussions, with six million monthly active users reported in March 2024 [Wodecki(2024)].
The popularity of these services, particularly among young adults, speaks to their growing acceptance as sources of social interaction and emotional support [Koulouri et al.(2022), Xygkou et al.(2024), Xygkou et al.(2023)]. Research suggests that artificial companions can reduce loneliness through various mechanisms, including direct companionship, acting as a catalyst for social interaction, facilitating remote communication, and providing reminders for social activities [Gasteiger et al.(2021)].
The perceived “realness” of the connection with AI companions is a crucial factor in their effectiveness. Users who perceive chatbots as having human-like qualities, such as consciousness or agency, often experience improved social health and reduced loneliness [Xia et al.(2024)]. This anthropomorphism can enhance the effectiveness of companion chatbots in mitigating feelings of isolation, though it carries a risk of leading to the user dehumanizing other people [Zehnder et al.(2021)]. Pataranutaporn & Liu et al. [Pataranutaporn et al.(2023)] also found that mental models altered experience – priming users with different beliefs about an AI’s motives (caring, manipulative, or neutral) significantly affected their assessment of the AI’s trustworthiness, empathy, and effectiveness. This suggests that users’ mental models of AI systems play a crucial role in shaping their interactions and experiences.
Previous work has suggested that the ability to empathize is inversely correlated with loneliness over a lifetime [Beadle et al.(2012)], but the question remains of what can be the role of a chatbot that simply performs empathy. Role-play has been shown to increase empathy [Bearman et al.(2015), Dewi et al.(2019)], and the potential of AI-facilitated role-play to influence empathy and, by extension, loneliness, represents an unexplored area in the research literature, despite the large number of users engaging in such interactions.
Existing research on the influence of artificial agents on socialization with other people has produced mixed results, with the work mostly qualitative in nature [Van der Loos(2014), Turkle(2011)]. One study suggested that artificial agents may end up replacing humans for task-oriented interactions, while in social contexts they are more likely to complement human social interactions [Xia et al.(2024)].
2.4 Human-AI Interaction for Emotional Support
Recent studies suggest that even limited amounts of AI companionship reduce the experience of loneliness [De Freitas et al.(2024)]. However, the long-term effects and potential risks of relying on AI for emotional support remain subjects of ongoing investigation and ethical debate [Li et al.(2023), Abd-Alrazaq et al.(2020)].
Skjuve et al. described how human-chatbot relationships (HCRs) evolve from an initial superficial phase through more affective exploration – with greater trust and self-disclosure – to a more stable state of emotional connection [Skjuve et al.(2021)]. This relationship development enables companion chatbots to provide various forms of social support, including companionship, emotional support, and informational guidance in a “safe space” that is always available but not physical [Ta et al.(2020)]. The availability and perceived safety of these interactions make them an attractive alternative for social support for those who may lack support from family and friends, such as LGBTQ individuals [Ma et al.(2024)]. The motivations behind these interactions are diverse, ranging from curiosity to explicit seeking of emotional support, with users engaging in topics from casual conversation to deep personal issues. These findings suggest that while chatbots can serve as valuable sources of emotional support, the nature and impact of these interactions vary significantly based on individual user characteristics and needs [Ta-Johnson et al.(2022)]. Though limited to one service, Siemon et al.’s analysis of Replika reviews [Siemon et al.(2022)] showed that users of Replika often use the service to cope with their loneliness, and their long-term interactions with Replika can be joyful and beneficial.
Recent studies by Jo et al. [Jo et al.(2023), Jo et al.(2024)] explored the deployment of an LLM-driven chatbot, CareCall, for public health interventions targeting socially isolated individuals. Their findings suggest that chatbots can alleviate public health workloads and mitigate loneliness, with users perceiving the system as personal and emotionally supportive. However, challenges such as trust, privacy, and the limitations of LLM-driven chatbots in addressing complex health needs were also highlighted in this study, emphasizing the importance of careful system design and managing stakeholder expectations.
Age can also affect how users interact with AI. Neves et al. [Neves et al.(2015)] examined the difficulties faced by older adults in adopting new communication technologies, highlighting the importance of user-centered design and support systems in facilitating technology acceptance. These insights are crucial when considering the potential of AI chatbots as a scalable solution for emotional support and loneliness reduction.
While companion chatbots offer promising avenues for emotional support, there are significant ethical considerations and potential risks associated with their use. These include the possibility of over-reliance on AI, consequences for mood and affect, delays in seeking professional help, and further withdrawal from human socialization [Pataranutaporn et al.(2021)]. A recent case of a teen suicide after extensive use of Character.ai raised concerns regarding whether or not the chatbot could be blamed for his death [Roose(2024)]; this case underscores the urgent need to understand the influence of companion chatbots on psychosocial well-being. The development of dependence on AI agents and the potential replacement of real human relationships raise important questions about the long-term societal impacts of widespread AI companion use.
More specific, granular information about users and their needs is needed. Our research seeks to identify the characteristics of users who benefit from chatbot interactions and users who may experience negative consequences. Our intent is to inform more targeted and ethical design practices in the rapidly evolving landscape of human-AI interaction towards the goal of positive computing in human-AI interaction – designing technology that supports psychological well-being [Calvo and Peters(2017)].
3 Methods
We surveyed companion chatbot users recruited from CloudResearch, administering several psychological scales and measures of chatbot usage frequency and session length. This study was approved by the institution’s IRB. Using the Qualtrics sample size estimator, we roughly estimated a minimum size of 385 with a 95% confidence level and a 5% margin of error. Our criteria for participants were that they must be:
  • 18 years old or older
  • fluent in English
  • regular users of companion chatbots (using them at least once a week for at least a month).
Participants were first recruited with a screening survey, only accessible on CloudResearch if they were fluent in English and had prior experience using AI. A researcher manually read through respondents’ brief free responses regarding their experience with companion chatbots to remove participants who used chatbots only for practical purposes (i.e. assisting with coding or interview preparation) or those who did not describe any personal experiences (i.e. only describing the service). Participants were screened and recruited through multiple rounds until we reached our target pool of qualified individuals. All qualifying participants were then invited to complete the main survey. We intentionally over-recruited to account for incomplete responses. After filtering out incomplete responses from the main survey, the final sample consisted of 404 valid responses (n = 404), slightly exceeding our initial target of 400.
We conducted the following analyses:
  • Participant Demographics: We calculated descriptive statistics to create a general overview of our participants.
  • Correlation Analysis: We calculated a Spearman correlation matrix with Benjamini-Hochberg correction as a starting point for exploring relationships in our numerical variables.
  • Multiple Regression Analysis: We performed an exploratory multiple regression analysis to construct a more nuanced model that can capture relationships between variables not reflected in the correlation analysis.
  • Participant Cluster Analysis: We performed K-means clustering of our numerical data to find distinct groups of users in our data.
  • Chatbot Usage Characteristics: We calculated descriptive statistics on multiple-choice questions regarding companion chatbot usage and examined participants’ free responses, both overall and grouped by the clusters we found. Our intention was to characterize our participants overall and provide more specific characteristics and examples of responses for each cluster. We also conducted thematic analysis of participants’ reasons for initial interest, motivations for continuing usage, and topics of conversation using a stratified random sample of 15 participants per cluster (total N = 105).
These analytical approaches were designed to address our primary research questions:
  1. 1.
    To explore motivations and uses of companion chatbots (RQ1), we analyzed the descriptive statistics of usage characteristics and examined participants’ free responses.
  2. 2.
    To investigate the relationship between chatbot usage and loneliness (RQ2), we employed correlation analysis and multiple regression analysis. These methods allowed us to identify potential mediating or moderating factors in this relationship.
  3. 3.
    To identify and characterize distinct user profiles (RQ3), we conducted K-means cluster analysis on our numerical data, including usage patterns, loneliness levels, and other psychological and social factors. We then examined chatbot usage characteristics within each cluster to provide more nuanced profiles.
Figure 3: A flowchart detailing the multi-method approach employed in this study. The diagram outlines the sequence of analytical steps, including participant recruitment, survey administration, data processing, correlation analysis, multiple regression modeling, cluster analysis, and the interpretation of both quantitative results and free-text responses.
We summarize these methods in Figure 3. By combining these quantitative and qualitative approaches, we aimed to provide a comprehensive understanding of companion chatbot usage patterns, their relationship with psychological factors such as loneliness, and the characteristics of different user groups.
3.1 Survey Design
Our survey instrument was designed to investigate the relationship between chatbot usage and loneliness, as well as potential mediating factors. The survey comprised both established scales (some adapted) and custom questions developed for this study. The full survey can be found as a Word document in Supplementary Material.
3.1.1 Established Scales and Adaptations
Our scale selection was guided by two principles: (1) measuring our primary variable of interest, loneliness, with a well-validated instrument, and (2) exploring potential mediating and moderating factors through an exploratory approach. While we acknowledge that including many scales risks multicollinearity, we addressed this through statistical controls in our analysis, detailed in the Analysis section.
Our primary dependent variable measure was the UCLA Loneliness Scale (ULS-8) [Hays and DiMatteo(1987)], demonstrating high internal consistency ($\alpha$ = .84) and construct validity through correlations with depression and inverse correlations with satisfaction with life, and social support [Wiseman et al.(1995), Wu and Yao(2008)]. It was selected for its brevity while maintaining psychometric properties of the original scale.
For our exploratory investigation of potential mediating and moderating factors, we included:
  • Lubben Social Network Scale (LSNS-6) [Lubben et al.(2006)]: Measures social isolation with strong internal consistency ($\alpha$ = .83). Selected to provide an objective measure of social connection to complement the subjective measure of loneliness.
  • Multidimensional Scale of Perceived Social Support (MSPSS) [Zimet et al.(1990)]: Assesses perceived social support. Selected for strong psychometric properties ($\alpha$ = .88) and wide usage [Wang et al.(2021)], as well as the known correlation between loneliness and social support [Ozbay et al.(2007), Wang et al.(2018), Schutter et al.(2021)].
  • Brief Rosenberg Self-Esteem Scale (B-RSES) [Monteiro et al.(2022)]: Evaluates self-esteem, a known risk factor leading to and increased by loneliness [Vanhalst et al.(2013)]. The brief scale was chosen for its shorter length but good internal consistency ($\alpha$ = .89 in a general population) and strong validity comparable to the original scale [Monteiro et al.(2022)].
  • Big Five Inventory (BFI-10) [Rammstedt and John(2007)]: Evaluates personality traits based on five factors: extraversion, neuroticism (tendency to towards anxiety and other negative feelings), agreeableness (tendency to be compassionate and cooperative), conscientiousness (tendency to be self-disciplined and goal-directed), and openness to experience (tendency to be open-minded, imaginative, and insightful) [McCrae and Costa(1999)]. It correlates highly with the full BFI (mean correlation .83) [Rammstedt and John(2007)]. Selected based on previous research linking personality traits to loneliness [Buecker et al.(2020)].
  • Attitudes Towards AI Scale [Schepman and Rodway(2020)]: Measures attitudes towards artificial intelligence with strong internal consistency for both the positive ($\alpha$ = .88) and negative ($\alpha$ = .83) subscales. Used the 8-item adaptation of the 16-item positive subscale, found in [Pataranutaporn et al.(2023)].
  • Human-Chatbot Interaction Effect Scale [Xia et al.(2024)]: Measures the effect of human-chatbot interaction on human-human interaction. We included measures of the respondent’s expectations for chatbot services ($\alpha$ = .85), trust in interactions with social chatbots ($\alpha$ = .86), perceived value of substituting people with chatbots ($\alpha$ = .91), trust in interactions with people ($\alpha$ = .93), and motivation and willingness to interact with people ($\alpha$ = .94). Changed phrasing to refer to only chatbots and people rather than service providers, shortened subscales based on factor loadings, and made two versions of the interpersonal communication intention subscales, one for familiar people and one for unfamiliar people.
  • Generalized Problematic Internet Use Scale (GPIUS2) [Caplan(2010)]: A validated measure of problematic internet use with strong psychometric properties ($\alpha$ = .91 for full scale, CFI = .95) [Caplan(2010)]. We adapted six items from the scale to specifically reference chatbot use, maintaining representation of subscales measuring compulsive use, cognitive preoccupation, and negative outcomes. Existing work shows that loneliness is both a cause and effect of problematic internet use [Kim et al.(2009)], and the GPIUS2 was developed with online social interactions in mind, making it well-suited for examining potential compulsive or problematic patterns in human-AI interaction.
  • State Empathy Scale [Shen(2010)]: A validated measure capturing affective, cognitive, and associative dimensions of empathy. The scale demonstrates strong construct validity and reliability across multiple studies, with confirmatory factor analyses supporting its unidimensional structure [Shen(2010)]. Adjusted wording to be more general for application to both chatbots and people. Selected as a way to assess how participants interpersonally connect with chatbots, as research suggests people process parasocial relationships similarly to interpersonal ones [Schiappa et al.(2005)] and there may be parallels between parasocial relationships and human-chatbot relationships.
  • Interpersonal Attraction Scale [McCroskey and McCain(1974)]: Measures three dimensions of attraction: social (desire for friendship), physical (appearance), and task (competence-based). The scale’s application to mediated relationships is supported by extensive parasocial interaction research [Schiappa et al.(2005)]. We shortened each subscale to four items based on factor loadings and created two versions: one for chatbots and one for close persons (excluding physical attraction for the latter due to potential family relationships).
  • Perceived Homophily Scale [Mccroskey et al.(1975)]: Measures perceived similarity in communication, which research shows is central to both interpersonal and parasocial relationship development [Schiappa et al.(2005)]. Modified to use unidirectional Likert statements (e.g., "like me" instead of "unlike me – like me"). Selected 6 items based on factor loadings and created parallel versions for chatbots and a close person.
  • Attributional Confidence Scale [Gudykunst and Nishida(1986)]: Assesses ability to predict others’ attitudes, feelings, and behavior [Clatterbuck(1979)]. Its relevance to human-AI interaction is supported by uncertainty reduction theory’s application to parasocial relationships; that is, relationships develop as people increase their ability to predict others’ behavior [Perse and Rubin(1989)]. Shortened to 5 items, removing items irrelevant to chatbots.
  • System Usability Scale (SUS) [Brooke(1996)]: A widely-validated measure of perceived system usability demonstrating high reliability ($\alpha$ = .91) across diverse interfaces [Bangor et al.(2008)]. Adjusted phrasing to refer specifically to chatbots. Chosen for consideration of potential links between usability and perceived closeness.
We acknowledge that modifying established scales may affect their validated psychometric properties. To manage survey length while maintaining measurement integrity, we shortened scales based on factor loadings from previous studies when available. While this approach allowed us to explore multiple potential factors, we recognize this as a limitation and address it through statistical controls for multicollinearity in our analysis, Benjamini-Hochberg correction for multiple comparisons, and conservative interpretation of results from modified scales. This exploratory approach allows us to identify potential factors warranting further investigation while maintaining statistical rigor through appropriate controls.
3.1.2 Custom Survey Items
In addition to established scales, we developed custom items to capture specific aspects of chatbot usage. Each item except usage characteristics required participants to elaborate on their choices and share thoughts not captured in the multiple-choice options as a free response, as our provided options were not expected to be exhaustive.
  • Usage Characteristics: Multiple-choice questions for frequency of chatbot use (5 options ranging from daily to once a month or less plus options indicating variation) and average session length (6 options: <15 min, 15-30 min, 10 min-1h, 1h-2h, 2h-4h, 4h+). Broad options were chosen to reduce the risk of inaccuracies from recall bias.
  • Motivations and Continued Use: Multiple-choice questions with required free-text elaboration on why participants began using chatbots and why they continue to use them; the options can be seen in Figure 5 (left graphs). These options were chosen through discussion with multiple researchers and based on uses and gratifications theory [Katz et al.(1973), Tanta et al.(2014)].
  • Conversation Topics: Questions about the topics participants discuss with chatbots, with the options chosen from discussions with multiple researchers (see top right graph of Figure 5).
  • Human vs. Chatbot Preference: Several items for participants’ preferences for human or chatbot interaction in various contexts, with the above conversation topics presented as options in a matrix question.
  • Perceived Impact on Relationships: Six Likert scales assessing participants’ perceptions of how chatbot use has affected their relationships with people by leading to (1) less interaction, (2) changes in conversation topics, and (3) greater ease, (4) greater enjoyment, (5) less engagement, and (6) less enjoyment in talking to people. These were selected through discussion with multiple researchers to roughly gauge participants’ perceived relationship changes.
3.2 Correlation and Multiple Regression
Data processing and statistical analyses were performed using Python, with the statsmodels library for regression analyses and scikit-learn for data preprocessing. We used a significance level of $\alpha=0.05$ for all statistical tests, unless otherwise specified. Usage frequency and session length were converted to linear numerical scores, with multiple responses averaged. All numerical variables were standardized using StandardScaler to improve interpretability and reduce potential multicollinearity [Aiken and West(1991)].
We computed a Spearman correlation matrix [Cohen et al.(2003)] with Benjamini-Hochberg correction for multiple comparisons[Benjamini and Hochberg(1995)]. For the regression analysis, we focused on session length as the independent variable, as it showed a significant correlation with loneliness where frequency did not.
Potential mediators were identified using the corrected p-values (p < 0.05). Mediation analysis followed Baron and Kenny’s [Baron and Kenny(1986)] approach, with Sobel tests [Sobel(1982)] assessing mediation significance. We tested moderation effects by creating interaction terms and examining their significance in regression models [Aiken and West(1991)]. To address potential multicollinearity, we calculated Variance Inflation Factors (VIF), considering variables with VIF > 5 for elimination [Hair et al.(2010), O’brien(2007)]. After removing the state empathy subscales, which were redundant with the full scale, all predictors were below the threshold.
To build our regression model, we started with all identified mediators, moderators, and confounders [Greenland et al.(1999), VanderWeele and Shpitser(2013)], including socialization and neuroticism on theoretical grounds. While retaining age and sex as control variables, we used backward elimination to remove predictors with high p-values (generally p > 0.1) to make a more parsimonious and interpretable model [Cohen et al.(2003)]. We then explored adding interaction terms one at a time, retaining significant interactions. Categorical variables (area type, education, relationship status, political party, race, employment status) were tested using ANOVA but were not retained due to non-significance.
We assessed model fit using R-squared, F-statistic, and condition number. Model assumptions were checked using Q-Q plots, residuals vs. fitted plots, and influence plots [Cook and Weisberg(1982), Fox(2015)]. The summary of the final model from our exploration can be seen in Figure 13.
To examine the potential mediators of our final model, we conducted a parallel multiple mediation analysis. This analysis was performed using a bootstrapping approach with 5,000 resamples, as recommended by [Preacher and Hayes(2008)]. The analysis controlled for several covariates, including age, sex, and personality traits (BFI). Confidence intervals (95%) for the indirect effects were calculated using the percentile method.
3.3 Cluster Analysis
We employed K-means clustering analysis to identify distinct user profiles of chatbot users based on their psychosocial characteristics. These profiles were identified based on a set of features with significant correlations (p > 0.05, corr > 0.3) to key variables, as well as features relevant to chatbot usage including usage frequency and measures of empathy and attraction with a chatbot; the features were reduced to 10 principal components. After some exploration, we used Chatbot Session Length, Frequency, Loneliness (ULS-8), and Problematic Use of Chatbot (GPIUS2) for our key variables. The optimal number of clusters was determined using a combination of quantitative metrics and qualitative interpretation:
  1. 1.
    We calculated the Within-Cluster Sum of Squares (WCSS) for a range of cluster numbers (2-14) and plotted the results to visualize the elbow method [Kodinariya and Makwana(2013)].
  2. 2.
    We computed silhouette scores for the same range of clusters to assess cluster cohesion and separation [Kaufman and Rousseeuw(2009)].
  3. 3.
    We examined the interpretability and meaningfulness of the resulting clusters for different k values, considering our research questions and domain knowledge.
The silhouette scores increased slightly as the number of clusters increased from 3 to 7, dropping to a local minimum at 8 clusters. The silhouette score fluctuated without consistent increases and there was no clear bend in the elbow method, we ultimately chose a 7-cluster solution. This decision was also based on the interpretability and distinctiveness of the resulting user profiles, aligning with our objective to understand the characteristics of different groups of users.
The final clustering was performed using the K-means algorithm with k=7, implemented in Python using the scikit-learn library [Pedregosa et al.(2012)]. The resulting clusters were then characterized based on their centroids and distribution of key variables. To validate our clustering solution, we employed a multi-faceted approach:
  1. 1.
    We conducted Kruskal-Wallis tests to examine differences among clusters for key variables (Session Length, Loneliness, and Problematic Use), as the data violated normality assumptions [Kruskal and Wallis(1952)].
  2. 2.
    After confirming significant differences with Kruskal-Wallis tests, we performed Dunn’s post-hoc test with Bonferroni correction for multiple comparisons [Dunn(1964)] to identify specific significant differences between clusters on our key variables (session length, loneliness, problematic use, and frequency).
  3. 3.
    We calculated the average silhouette score to assess the quality of cluster separation [Rousseeuw(1987)].
  4. 4.
    We performed a Multivariate Analysis of Variance (MANOVA) to test for overall differences among clusters across multiple dependent variables simultaneously [Huberty and Olejnik(2007)].
  5. 5.
    We used Box’s M-test to check the assumption of homogeneity of covariance matrices for the MANOVA [Box(1949)].
These validation techniques provide a comprehensive assessment of the cluster solution’s statistical robustness and meaningful differentiation among user profiles.
3.4 Interpretation of Quantitative Results and Free-Text Responses
To provide a comprehensive understanding of our findings, we employed a two-pronged approach to interpreting the results of our quantitative analyses. First, we calculated descriptive statistics on companion chatbot usage characteristics for both our entire dataset and the clusters we identified through K-means analysis. This provided a quantitative overview of usage patterns and user profiles.
Second, we examined the free-text responses associated with multiple-choice items to complement our quantitative findings. We conducted a thematic analysis [Braun and Clarke(2006)] of three free-response items (participants’ reasons for initial interest, motivations for continuing usage, and topics of conversation) using a stratified random sample; 15 participants were randomly selected from each cluster while maintaining equal representation across clusters (total N = 105). A researcher read through the sample of free responses, developed a codebook, and extracted overall themes for each examined item. The researcher developed initial codes by identifying recurring patterns, then grouped these codes into broader themes that captured key aspects of users’ reasons for engaging with companion chatbots.
By integrating these approaches, we were able to capture both broad trends identified through our quantitative analyses and individual nuances in user experiences expressed in their own words. This method provided a richer interpretation of the data, offering insights into the motivations, perceptions, and experiences of companion chatbot users that might not be fully captured by quantitative measures alone.
4 Results
4.1 RQ1: What motivates the use of companion chatbots, and what do people use them for?
Figure 4: Percentage of total responses for multiple-choice questions regarding chatbot usage patterns among participants. Top left: chatbot session length. Top right: chatbot usage frequency. Bottom left: chatbot usage timeframe. Bottom right: most used chatbot platforms.
Figure 5: Percentage of participant responses for multiple-choice questions regarding their use of companion chatbots. Top left: reasons for using companion chatbots. Top right: conversation topics with chatbots. Bottom left: motivations for continued use. Bottom right: preferred devices for companion chatbot use.
Figure 6: Codes and themes for initial and continued usage of companion chatbots from thematic analysis of a stratified random sample of participants (15 per cluster, total N = 105).
4.1.1 Motivations for initial interest and continued use of companion chatbots
Analysis of both quantitative and qualitative data revealed diverse motivations for chatbot engagement, with several key themes emerging across both initial adoption and continued use. Descriptions of these themes can be found in Figure 6.
Technology exploration emerged as the predominant motivation, manifesting quantitatively in curiosity about AI and chatbots (30.89% initial use, 20.91% continued use). Users expressed interest in testing the technology’s capabilities and limitations: “I wanted to see if it could actually emulate a real person.”
Recreation represented another major theme, encompassing both entertainment (26.35% initial, 23.69% continued) and creative activities (17.81% initial use). Users enjoyed role-playing with characters and leveraging chatbots for creative projects: “I use chatbots to practice conversations for my D&D game, creating NPC personalities and dialogue.”
The practical utility theme emerged particularly in continued use, with users valuing chatbots for task assistance and information seeking. As one user explained: “I usually use chatbots to help with ideas like dinner ideas or information I may want to find out without using a typical search engine.” This theme aligns closest with the quantitative findings showing passing time and reducing stress (12.88% continued use) as a notable factor, though there were no directly-related multiple-choice options originally.
Another theme not explicitly reflected in the initial multiple-choice options was the safe space theme, which emerged to encompass the value of chatbots as non-judgmental environments for self-expression an emotional support. One participant noted, “I sometime feel lonely and just want to be left alone, during this time I like chatting with my AI companion because I feel safe and won’t not [sic] be judged for the inadequate decisions I have made.”
Social motivations, while less prevalent, are significant for some users. These motivations encompassed themes companionship, self-improvement, and lifestyle integration. Initial use showed loneliness (12.24%), social skill development (7.24%), and relationship seeking (3.53%) as drivers, while continued use revealed ongoing needs for companionship (9.86%), emotional experiences (6.52%) and self-discovery (6.12%). Some users explicitly sought social improvement: ‘I’m autistic, so I’m not the best at social interaction, and I found that practicing with chatbots helped.” The lifestyle integration theme emerged specifically in continued use, with some users incorporating chatbots into their daily routines: “It has become part of my life and I want to continue to learn and evolve with new technology.” This suggests that for some users, chatbots transition from novelty to utility and integration over time.
Our thematic analysis enriches our understanding beyond pure statistics, revealing how different motivations interweave and evolve throughout users’ engagement with companion chatbots. While initial interest often stems from curiosity and entertainment, continued use appears more deeply rooted in practical and emotional benefits.
Figure 7: Codes and themes for topics of conversation with companion chatbots from thematic analysis of a stratified random sample of participants (15 per cluster, total N = 105).
4.1.2 Topics of conversation with companion chatbots
Analysis of both structured responses and free-text data revealed five major themes in how users engage with companion chatbots (see Figure 7).
Casual exchange, including casual conversations (26.28%) and entertainment (21.66%), represents the most common form of interaction. Participants used the chatbot to review past events (12.60%), such as talking about their day: “I often chat with my AI about random things like my day, and it feels like I’m having a normal conversation without the pressure of talking to a real person.”
Emotional disclosure emerged as a notable theme, encompassing mental health discussions and personal problems. This is reflected in the quantitative findings of personal issues and mental health (14.17%) and interpersonal issues and drama (10.96%) as notable topics. One participant responded, “I can talk about my problems, and it’s like having a private conversation with no fear of being criticized.”
Though less prevalent in the free responses, the theme of creative development manifested through topics of writing, roleplay, and creative ideation, with users leveraging chatbots as collaborators: “I typically ask for ideas for my marketing art! I do a lot of graphic design for a lot of different people and sometime run low on creative ideas.”
Knowledge seeking involved users turning to chatbots for information and advice, including discussion of future plans (11.70%). One individual mentioned, “I use the chatbot to ask about random things like historical facts or recommendations for books.” Others used companion chatbots as tools for specific tasks, such as recipe ideas.
A rare theme emerging from the free responses was intimate exchange, involving romantic and sexual interactions: “[Some]times I just want to feel pleasure and the interaction is very sexual. Other times I want to feel love and connection and it’s more romantic.”
We note that the themes identified in free responses sometimes diverge from the structured response options. This diverse range of conversation topics highlights the versatility of companion chatbots in meeting users’ needs for entertainment, emotional support, and intellectual stimulation.
4.2 RQ2: What is the relationship between companion chatbot usage and loneliness, and what factors mediate or moderate this relationship?
4.2.1 Correlations in psychological, social, and companion chatbot usage characteristics
For our prespecified analyses of usage vs. loneliness, we found a small but significant correlation between chatbot session length and loneliness ($r=0.101,p=0.042$), but not between loneliness and usage frequency ($r=0.053,p=0.287$), as shown in Figure 9.
Several other notable correlations emerged in our exploratory analysis using Benjamini-Hochberg corrected p-values. Chatbot usage (both length and frequency) showed moderate positive correlations with state empathy towards chatbots ($r>0.3,p<0.001$) and various forms of attraction towards chatbots ($r>0.18,p<0.001$). Loneliness demonstrated significant correlations with personality traits of extraversion, ($r=-0.188,p<0.001$), conscientiousness ($r=-0.335,p<0.001$), agreeableness ($r=-0.333,p<0.001$), and neuroticism ($r=0.526,p<0.001$), aligning with Buecker et al.’s [Buecker et al.(2020)] meta-analysis. Additionally, problematic use of chatbots (GPIUS2) showed moderate correlations with both session length ($r=0.303,p<0.001$) and frequency ($r=0.337,p<0.001$). See Figure 8 for a correlation matrix of our continuous variables.
While these correlations provide initial insights into the relationships between variables, they represent associations rather than causation. A more detailed examination of these relationships through multiple regression analysis follows.
Figure 8: Correlation matrix of psychological, social, and usage characteristics. Variables are sorted into general categories, and names have been altered for readability. The original names can be seen in Figure 18 of Appendix A.
Figure 9: Plots of loneliness (ULS-8) and problematic usage (GPIUS2) vs. usage (frequency and session length). Top left: loneliness vs. chatbot usage frequency. Top right: loneliness vs. chatbot session length. Bottom left: problematic use vs. chatbot usage frequency. Bottom right: problematic use vs. chatbot session length..
Figure 10: The final model from our exploratory multiple regression, depicting the relationships between chatbot session length, loneliness, and related factors.
4.2.2 Multiple Regression Model
[…]
Figure 11: Interaction effects between session length and social attraction to a close person. Higher social attraction is associated with a stronger positive relationship between session length and loneliness. All variables are centered; SD = standard deviation.
A parallel multiple mediation analysis (Figure 12) revealed that problematic internet use (GPIUS2) significantly mediated the relationship between session length and loneliness ($\beta=0.0183$, 95% CI $[0.0027,0.0403]$), while social network size (LSNS) did not show significant mediation ($\beta=0.0007$, 95% CI $[-0.0304,0.0340]$). The total indirect effect was not statistically significant ($\beta=0.0191$, 95% CI $[-0.0149,0.0563]$).
Figure 12: Results of bootstrapped parallel multiple mediation analysis (5,000 resamples) examining the mediating effects of social network size (LSNS) and problematic chatbot use (GPIUS2) on the relationship between chatbot session length and loneliness, including lower and upper confidence intervals (CI_Lower and CI_Upper).
Figure 13: Model summary of an ordinary least squares regression for the effect of chatbot session length on loneliness (ULS-8), using standardized variables. Model Formula: Loneliness (ULS-8) $\sim$ Session Length Score + Social Attraction towards Chatbot + Problematic Use (GPIUS2) + Socialization (LSNS) + Social Attraction towards Close Person + Neuroticism + Agreeableness + Sex + Age + Session Length Score $\times$ Social Attraction towards Close Person
4.3 RQ3: Can we identify and characterize distinct user profiles based on usage patterns, loneliness levels, problematic use, and other psychological and social factors?
4.3.1 Overall characteristics
Participants ranged in age from 18 to 73, primarily 30-35. Most (93.6%) were from the United States, white/Caucasian (66.3%), and employed full-time (61.6%), with a broad distribution of household income. For relationship status, most were single (42.1%) or married (32.4%). 59.7% of participants identified as men and 36.6% identified as women. Additional demographic characteristics are in Figure 14.
Most participants utilize Snapchat MyAI (31.6%), Character.ai (30.4%), or Replika (22.6%) and tend to spend 15 and 30 minutes in each session with a companion chatbot (42.6%), with 35.4% reported sessions shorter than 15 minutes. Very few users spend extended periods chatting, with only 1.28% having sessions exceeding 4 hours. Most participants have been using their companion chatbots for a few months (43.9%) to around a year (35.2%), with only 7.7% of participants having used their chatbots for multiple years and 13.2% for around a month or less. More details on usage can be seen in Figure 4 and Figure 5.
Figure 14: Demographic table for study participants, including overall percentage for each category and the difference in percentage from the overall percentage for each cluster, with darker colors indicating a greater difference. Blue: cluster has a lower percentage of the category compared to overall. Red: cluster has a higher percentage of the category compared to overall.
4.3.2 Clusters of users of companion chatbots
Our K-means cluster analysis revealed seven distinct user profiles among companion chatbot users, each exhibiting unique behaviors, motivations, and psychological characteristics. The resulting clusters are visualized in the heatmap and dendrogram presented in Figure 15. The demographics of each cluster compared to overall demographics can be seen in Figure 14, with percentages found in Figure 20 of Appendix B.
Validation of the cluster solution using Kruskal-Wallis tests revealed significant differences among clusters in terms of session length, loneliness (ULS-8), problematic use of chatbots (GPIUS2), and frequency of use (all $p<0.001$). The average silhouette score ($-0.0413$) suggested some overlap between clusters, which is not uncommon in complex psychological data. A MANOVA test confirmed significant multivariate differences among the clusters when considering all variables simultaneously (Wilks’ $\lambda=0.2420,F(24,1375.71)=28.77,p<0.001$). All four MANOVA test statistics (Wilks’ lambda, Pillai’s trace, Hotelling-Lawley trace, and Roy’s greatest root) were significant ($p<0.001$), providing robust evidence for cluster differences across multiple dimensions. Box’s M-test indicated significant heterogeneity of covariance matrices ($\chi^{2}=181.96$, $p<0.001$), which does not invalidate the clustering but suggests the need for cautious interpretation. This heterogeneity may affect the reliability of post-hoc comparisons and could indicate that the relationship between variables differs across clusters. While these results provide strong support for the distinctiveness of the identified clusters, they should still be considered exploratory, providing a foundation for more targeted investigations in future research.
Figure 15: Right: Cluster means of standardized variables of chatbot usage (frequency and session length), loneliness, and problematic use of chatbots for our seven user clusters. Left: a dendrogram depicting the degree of similarity between each cluster.
Figure 16: Heatmap of cluster means for standardized variables in each cluster.
The seven identified clusters represent distinct patterns of chatbot engagement and psychological characteristics:
  • 0: Disengaged Light Users (7.67%): Minimal chatbot engagement with average loneliness levels. Show neutral attitudes toward AI technology and average social support levels. Primary usage is brief (<15 min) casual conversations, with main motivations being curiosity to start use and practical utility for continued use. Average age is 32; for gender, 65% are men and 35% are women.
  • 1: Well-Adjusted Moderate Users (23.02%): Average usage with low loneliness. Demonstrate high extraversion, low neuroticism, and strong social networks. Show positive attitudes toward both AI and human interactions. Main motivations of curiosity to start use and emotional support to continue use. Average age is 36; 67% men, 33% women.
  • 2: AI-Wary Light Users (11.88%): Average to low usage with low loneliness. Show high skepticism toward AI technology while maintaining strong human connections. Users were interested in chatbots for practical utility and recreation, with continued use motivated by practical purposes, such as seeking knowledge and advice. Average age is 34; 65% men, 35% women.
  • 3: Lonely Moderate Users (13.61%): Average usage with high loneliness. Show low socialization but positive attitudes toward chatbots. Notable for seeking emotional support and companionship through chatbot interactions, with some participants integrating chatbots into their lifestyle. Average age is 35; 51% men, 45% women, 4% other.
  • 4: Fulfilled Dependent Users (18.32%): High usage with below-average loneliness. Show strong emotional engagement with chatbots while maintaining adequate social connections. Report positive impacts on real-world social interactions. Average age is 35; 61% men, 37% women, 1% other.
  • 5: Lonely Light Users (14.60%): Average usage with high loneliness. Demonstrate high neuroticism and low self-esteem. Users seek out chatbots for emotional support and companionship, often discussing personal issues and sometimes engaging in intimate relationships with chatbots. Average age is 34; 47% men, 52% women.
  • 6: Socially Challenged Frequent Users (10.89%): Average frequency but short sessions, with high loneliness. Show low extraversion and social support, with notably low trust in both familiar and unfamiliar people. Both initial and continued use is motivated by a desire for companionship, with a few rare cases engaging in intimate relationships with their chatbot. Casual exchange was a less common theme of conversation. Average age is 36; 61% men, 39% women.
Detailed heatmaps including all numerical scales can be seen in Figure 16. A version with the original values of the scales can be found in Figure 19 along with detailed cluster descriptions in Appendix B. The patterns of themes for each cluster’s motivations for use and topics of discussion can be seen in Figure 6 and Figure 7.
Post-hoc analysis using Dunn’s test with Bonferroni correction revealed distinct patterns among clusters across our key variables (Figure 17). For session length, Fulfilled Dependent Users (Cluster 4) showed significantly longer sessions than all other clusters (p < .001) except Well-Adjusted Moderate Users. Disengaged Light Users (Cluster 0) had significantly shorter sessions than most other clusters (p < .001).
Usage frequency showed less variation between clusters, with only Disengaged Light Users showing significantly lower frequency than all other clusters (p < .001), and Fulfilled Dependent Users showing higher frequency than AI-Wary Light Users and Lonely Light Users (p < .01).
For loneliness (ULS-8), three clusters - Lonely Moderate Users, Lonely Light Users, and Socially Challenged Frequent Users (Clusters 3, 5, and 6) - showed significantly higher levels than Well-Adjusted Moderate Users and AI-Wary Light Users (p < .001). Fulfilled Dependent Users maintained significantly lower loneliness levels than these high-loneliness clusters (p < .001).
Problematic use (GPIUS2) was highest in Lonely Moderate Users and Fulfilled Dependent Users (Clusters 3 and 4), who showed significantly higher levels than all other clusters (p < .001) except each other. AI-Wary Light Users (Cluster 2) demonstrated the lowest levels of problematic use (p < .001).
Figure 17: Significant pairwise differences between clusters on key variables based on Dunn’s test with Bonferroni correction. Cell background color indicates significance of differences (darker green = presence of greater significance in differences), with text color indicating direction (red = greater than, blue = less than). For example, >2*** indicates significantly greater values than Cluster 2 at p < .001. Significance levels: *: p < .05, **: p < .01, ***: p < .001.
5 Discussion
Our study provides a detailed examination of the relationship between companion chatbot usage and loneliness, offering some insights into the complex interplay of psychological, social, and technological factors that shape this relationship.
5.1 Understanding the Relationship Between Companion Chatbot Usage and Loneliness
To answer RQ2, “What is the relationship between companion chatbot usage and loneliness, and what factors mediate or moderate this relationship?” we look to our regression analysis. While session length showed no significant direct relationship with loneliness, our regression analysis revealed a more nuanced picture through mediating and moderating factors. Problematic use emerged as a significant mediator, suggesting that increased usage may create more opportunities for problematic behaviors to develop. This aligns with existing research connecting problematic internet use as both a cause and effect of loneliness [Kim et al.(2009)], indicating a similar pattern may exist for companion chatbot use.
The negative association between usage and social network size warrants attention, though causality remains unclear. Users might reduce human interactions as they spend more time with chatbots, or they might turn to chatbots to fill an existing social void. The positive relationship between usage and social attraction towards chatbots is logically consistent – more interaction time allows deeper attachments to form. However, whether this attraction leads to loneliness through unfulfilled relational needs or loneliness enhances attraction to chatbots is uncertain.
Among confounding variables, neuroticism showed a strong positive association with loneliness, while agreeableness demonstrated a strong negative association, aligning with previous research on personality traits and loneliness [Buecker et al.(2020)], reinforcing the importance of controlling for these factors in studies of technology use and social well-being.
An interesting finding was the positive interaction term between session length and social attraction towards a close person. This suggests a nuanced relationship between chatbot usage and loneliness for socially attracted individuals. This could be interpreted as a mismatch between social needs and chosen social interactions, where time spent with chatbots may be perceived as time not spent with valued human connections. We can also consider a reversed causality: perhaps a very lonely user becomes more socially attracted to people as they spend more time with their chatbot, or non-lonely users could be satisfied with their chatbots and subsequently become less socially attracted to people.
5.2 Understanding Clusters of Chatbot Users
We shall first discuss the overall characteristics of our participants, which we explored for RQ1, “What motivates the use of companion chatbots, and what do people use them for?” Our analysis revealed that despite companion chatbots’ marketing as social tools, most participants primarily used them for technological exploration and entertainment rather than friendship or intimate relationships. This usage pattern may parallel other solitary activities like gaming or reading, which can affect socialization and lead to overindulgence. However, our cluster analysis revealed distinct user profiles demonstrating more complex relationships with these technologies. This brings us to RQ3, “Can we identify and characterize distinct user profiles based on usage patterns, loneliness levels, problematic use, and other psychological and social factors?”
The most striking contrast emerges between Socially Fulfilled Dependent Users (Cluster 4) and Lonely Moderate Users (Cluster 3). Both exhibit high problematic use, yet their outcomes differ dramatically. Socially Fulfilled Dependent Users maintain high well-being despite (or perhaps because of) their extensive chatbot engagement. Their high empathy and attraction towards chatbots, combined with positive reported effects on real-world social interactions, suggests that intensive chatbot use can coexist with healthy social functioning. These findings parallel work by Snodgrass et al. showing that more intensively involved gamers, those exhibiting what appears to be “addictive” play, experience the greatest psychosocial benefits from their gaming – including reduced loneliness [Snodgrass et al.(2018)]. However, unlike gaming where more intense involvement tends to predict better social outcomes, we found that Lonely Moderate Users with similar usage patterns experienced markedly different results. For chatbots, the nature of engagement may matter more than intensity alone. This challenges concerns about the negative effects of AI replacing human relationships [Pataranutaporn et al.(2021), Turkle(2011)] and merits further investigation into how these users maintain this balance.
Conversely, Lonely Moderate Users demonstrate how similar usage patterns can lead to markedly different outcomes. Their high loneliness, low socialization, and preference for chatbots over humans raise concerns about whether chatbot use exacerbates social withdrawal or merely reflects existing social difficulties. Their pattern of emotional disclosure to chatbots suggests these tools may serve as their primary outlet for emotional expression. We suggest further investigation of whether chatbots have a positive or negative effect on these individuals.
Socially Challenged Frequent Users (Cluster 6) and Lonely Light Users (Cluster 5) are another interesting pair to consider. Socially Challenged Frequent Users are most similar to Lonely Moderate Users (3), though they have average problematic use and considerably below-average measures of trust and empathy, even for those they feel close to. These individuals likely need intervention beyond what a companion chatbot can currently provide, yet companion chatbots may be the only entities they are willing to engage with. Both Socially Challenged Frequent Users and Lonely Light Users seek out chatbots for companionship, and it is largely only in these two groups that participants discussed having or seeking a romantic or sexual relationship with their chatbot. The participants in these two clusters most strongly reflect the Media Equation, in that their responses to chatbots may be similar to their social responses to people [Reeves and Nass(1998)]. Of concern is whether or not Lonely Light Users might transition to Socially Challenged Frequent Users over time, losing trust and empathy with people as they engage more with chatbots. Considering that loneliness is a risk factor for problematic use [Kim et al.(2009)], understanding how these users’ characteristics can change over time, and how chatbot systems can detect these at-risk individuals and intervene, may be valuable for future research.
Well-Adjusted Moderate Users (Cluster 1), the largest group, demonstrate a potentially optimal engagement pattern. With moderate usage and lower emotional connection to AI, they maintain strong human connections while leveraging chatbots primarily as tools for emotional processing rather than social companions. This suggests that chatbots might effectively reduce loneliness without necessarily functioning as relationship substitutes. Many disclose their emotional issues to chatbots, valuing the safe space for expression they provide, with a few considering the chatbot as part of their lifestyle. Chatbots might effectively reduce loneliness without necessarily functioning as relationship substitutes. Further investigation into causality and mechanisms of psychosocial changes may prove fruitful for understanding how companion chatbots can be designed to reduce loneliness, besides the typical strategy of pushing for greater emotional intimacy with the user. For consideration, recent work has shown that social support acts as a significant mediator between psychological problems and problematic gaming [Malak et al.(2023)], which may help explain how Cluster 4, with poor perceived social support, exhibits high loneliness – as opposed to Cluster 1, with high perceived social support and low loneliness – when both have moderate chatbot usage.
The AI-Wary Light Users (Cluster 2) provide an important counterpoint, showing that emotional distance from chatbots can coexist with low loneliness. These users show low emotional attachment to chatbots and low opinions of AI in general. This group uses chatbots mostly for practical purposes and recreation, perhaps without seeking or developing emotional connections. Many traits associated with loneliness – neuroticism, agreeableness, socialization, and self-esteem) are fairly average, but their loneliness is the second lowest of all clusters. Considering this in comparison to other clusters suggests that emotional investment in AI companions may not be necessary or even optimal for all users.
These diverse profiles highlight that the impact of companion chatbot use on loneliness and social well-being varies significantly based on individual characteristics and usage patterns. This heterogeneity suggests the need for personalized approaches in both chatbot design and potential intervention strategies.
5.3 Implications and Future Directions
Our findings extend beyond simple social displacement theories [Kraut et al.(1998)]suggesting a framework where companion chatbots serve different social functions based on user characteristics and usage patterns. This aligns with and extends the Parasocial Contact Hypothesis [Schiappa et al.(2005)], which suggests that media characters can provide meaningful social experiences. Chatbot interactions may similarly provide parasocial benefits, but with important distinctions based on user traits and usage patterns.
The mediating role of problematic use in the relationship between usage and loneliness contributes to theories of technology addiction and compensatory internet use [Kardefelt-Winther(2014)]. Excessive chatbot use may reflect attempts to compensate for social deficits, similar to problematic social media use, but with unique characteristics due to the interactive nature of companion chatbots. This extends existing frameworks for understanding problematic technology use to include AI companionship specifically.
5.3.1 Theoretical Contributions
Potential Benefits and Risks. For some individuals, particularly those with strong existing social networks and positive psychological profiles, companion chatbots may serve as a beneficial supplement to human social interaction. These users may experience higher social confidence in parallel with their chatbot use. Our data suggest that neurotic individuals, who tend to be lonely and socially isolated, may experience reduced loneliness when spending more time with chatbots, but previous research suggests that more socially isolated individuals are at greater risk of problematic use of online services [Reer et al.(2019), O’Day and Heimberg(2021)]. As our model should be interpreted cautiously, the risk of chatbots replacing human interactions and potentially exacerbating loneliness and social withdrawal should be considered.
Implications for HCI Design and Practice. Our findings on problematic use suggest the need for built-in mechanisms to detect patterns of potentially harmful usage, provide user feedback about usage patterns, and implement interventions when problematic patterns emerge. Additionally, offering resources for maintaining healthy human relationships alongside chatbot use may prove beneficial if they can help improve the user’s social support.
Further, if the user’s traits can be collected, designing adaptive interfaces for social chatbots based on user traits to encourage or discourage certain uses may be an avenue for mitigating risk of negative effects for at-risk individuals (such as those with high neuroticism and signs of problematic use). Chatbots could potentially be tailored to different user types, with features and interaction styles that complement rather than replace human social interaction for vulnerable users.
For individuals with limited access to human social support, companion chatbots could serve as an intervention to reduce loneliness and its associated health risks. However, the design and implementation of these interventions must be carefully considered. Future HCI research should consider multiple pathways of influence when studying and designing AI companion technologies.
Ethical Considerations & Long-term Societal Impacts. The potential for some users to engage in deep, emotional relationships with chatbots raises ethical concerns. Developers and policymakers need to consider how to balance the potential benefits of AI companionship with the risks of over-reliance on artificial relationships. This may involve implementing safeguards or guidance systems within chatbots to encourage healthy usage patterns and human social interaction.
As AI technology advances and companion chatbots become more sophisticated, it is crucial to consider the long-term societal impacts of widespread AI companionship. Questions arise about how these technologies might shape social norms, expectations for human relationships, and the development of social skills, particularly for younger generations. Should we as a society push for less artificial companionship, or should we encourage a healthy mix of artificial and real companionship? Consider the case of a profoundly lonely and emotionally vulnerable individual. They might cling to a relationship that can provide any crumb of affection, no matter how unhealthy it becomes. Could companion chatbots fulfill that desperate need in a beneficial way, even if “unhealthy attachment” to the chatbot is part of the process? Perhaps alleviating loneliness enough to reduce the individual’s vulnerability could allow them to become more conscientious about seeking out and maintaining healthy, respectful human relationships.
5.4 Future Research Directions
Several key areas warrant further investigation. Longitudinal studies would be particularly valuable in understanding the causal relationships between chatbot usage, loneliness, and social outcomes over time. Such studies could track changes in social behaviors, relationship patterns, and psychological well-being as individuals engage with companion chatbots over extended periods.
Experimental studies manipulating specific aspects of chatbot design or usage could help identify features that promote healthy engagement while minimizing risks of problematic use or social displacement. These studies might explore factors such as conversation topics, interaction frequency, or anthropomorphic qualities of chatbots and their impacts on user experiences and outcomes.
Given the complex interactions observed in our model, future research should also investigate potential non-linear relationships between chatbot usage and psychological outcomes. This could involve more sophisticated statistical modeling techniques or qualitative approaches to capture nuanced patterns of usage and their effects.
The cultural context in which companion chatbots are used likely plays a significant role in shaping their impact. Cross-cultural studies could explore how different societal norms, communication styles, and attitudes towards technology influence the relationship between chatbot usage and loneliness.
Focused studies on specific user groups, such as the clusters we identified or specific demographic groups such as adolescents, older adults, or individuals with social anxiety, could yield valuable insights for developing targeted designs or guidelines.
As companion chatbots become more prevalent, it will be crucial to investigate their long-term effects on social skills, relationship expectations, and overall social functioning.
5.4.1 Limitations and Considerations
Survey design and recruitment. As our survey consists entirely of self-reported measures for usage characteristics and psychosocial measures, we face a considerable risk of recall bias. Our custom instruments for usage characteristics offered fairly broad options to reduce potential inaccuracies, though this also makes them coarse measures. Conducting a study involving monitoring or reporting of actual use would provide more specific and accurate usage data, and is in our plans for future work.
While our adaptations of established scales were based on factor loadings from previous studies and tailored to our research context, we acknowledge that these modifications may affect the scales’ validated psychometric properties. Future work should consider validating these adapted scales specifically for chatbot-related research. Due to the extensive nature of our survey, we prioritized reducing participant fatigue over maintaining the full length of each scale. This trade-off allowed us to explore a broader range of factors but may have impacted the depth of measurement for individual constructs. Additionally, our participants were primarily white/Caucasian individuals from the United States; applying our findings to other demographics should be done cautiously.
Clustering. We note that our clusters have their own set of limitations. While the number of clusters has statistical grounding, the final number can be considered somewhat arbitrary and based on our subjective judgments of interpretability. The heterogeneity of covariance matrices indicated by Box’s M-test requires caution in interpreting these results. Even so, this 7-cluster solution provides a starting point for further work on styles of use.
Regression Model. The exploratory nature of our model selection process warrants cautious interpretations and further investigation into the relationships it implies. The assumption of linear relationships may not capture complex nonlinear relationships in the population of users of companion chatbots. Further, due to the numerous predictors involved in the exploration and the final model, the risk of Type II errors is considerable [Lavery et al.(2019)]; our analysis may not capture all potential relationships of these variables.
Additionally, while we can theorize about the directions of relationships between variables, our cross-sectional design limits causal inferences. Future research employing experimental designs or longitudinal studies could provide more definitive evidence of causal relationships and explore potential non-linear effects. We plan to follow up with our participants to see changes in psychological, social, and usage characteristics at different time points.
6 Conclusion
This study examines the relationship between companion chatbot usage and loneliness, developing a model that explains approximately 50% of variance in loneliness through the complex interplay of psychological, social, and technological factors. Our findings reveal that while chatbot session length does not directly predict loneliness, this relationship is mediated by problematic use and moderated by social attraction to close others, highlighting the nuanced ways in which AI companions influence social well-being.
The model we developed through regression analysis identifies key mechanisms, including problematic use as a mediator and social attraction as a moderator, while accounting for personality traits such as neuroticism and social network characteristics. This suggests that the impact of companion chatbot usage on loneliness operates through complex pathways rather than direct effects.
Our cluster analysis further underscores this complexity, identifying seven distinct user profiles that demonstrate how similar usage patterns can lead to markedly different outcomes. For instance, socially fulfilled dependent users maintain healthy relationships despite intensive chatbot use, while lonely moderate users with similar usage patterns show signs of social withdrawal. These contrasting outcomes emphasize the importance of individual characteristics and social contexts in shaping the effects of AI companionship.
As AI companions become increasingly prevalent, these findings have crucial implications for design and deployment. The identification of distinct user profiles and key mediating factors suggests the need for personalized approaches and built-in mechanisms to detect and intervene in potentially harmful usage patterns. The contrasting outcomes between different user types raise important questions about how to design AI companions that support psychosocial well-being.
Future research should explore the causal relationships and long-term impacts of AI companionship through longitudinal studies while investigating how different design approaches might better support healthy engagement patterns. As we continue to navigate the evolving landscape of human-AI interaction in an increasingly digital world, understanding these complex relationships becomes crucial for developing technologies that genuinely enhance human well-being.
7 Acknowledgments
Statistical support was provided by data science specialist Jinjie Liu, at the Institute for Quantitative Social Science, Harvard University.
Appendix A Significant Correlations with Key Variables
Below, we include a list of significant (corrected $p<0.05$) correlations with chatbot usage frequency score, chatbot session length score, and loneliness (ULS-8). Negative correlations have text colored in blue. Additionally, the correlation matrix with original variable names is seen in Figure 18.
Figure 18: Correlation matrix of psychological, social, and usage characteristics with original variable names.
Significant correlations with Chatbot Usage Frequency:
  • Chatbot Session Length: $r=0.3497,p<0.001$
  • Problematic Use of Chatbot (GPIUS-2): $r=0.3374,p<0.001$
  • System Usability (SUS): $r=0.2457,p<0.001$
  • Conscientiousness (BFI-10): $r=0.1511,p0.005$
  • Attitudes towards AI: $r=0.2976,p<0.001$
  • State Empathy towards Chatbot: $r=0.3207,p<0.001$
  • Affective State Empathy towards Chatbot: $r=0.2472,p<0.001$
  • Cognitive State Empathy towards Chatbot: $r=0.2805,p<0.001$
  • Associative State Empathy towards Chatbot: $r=0.3416,p<0.001$
  • Social Attraction towards Chatbot: $r=0.2238,p<0.001$
  • Physical Attraction towards Chatbot: $r=0.1856,p<0.001$
  • Task Attraction towards Chatbot: $r=0.1652,p0.002$
  • Perceived Similarity with Chatbot: $r=0.2525,p<0.001$
  • Expectations towards Chatbot (HCI/HHI): $r=0.2560,p<0.001$
  • Trust towards Chatbot (HCI/HHI): $r=0.2931,p<0.001$
  • Preference for Chatbot over Humans (HCI/HHI): $r=0.2084,p<0.001$
Significant correlations with Chatbot Session Length:
  • Chatbot Usage Frequency: $r=0.3497,p<0.001$
  • Problematic Use of Chatbot (GPIUS-2): $r=0.2549,p<0.001$
  • System Usability (SUS): $r=0.1962,p<0.001$
  • Attitudes towards AI: $r=0.2331,p<0.001$
  • State Empathy towards Chatbot: $r=0.3016,p<0.001$
  • Affective State Empathy towards Chatbot: $r=0.2475,p<0.001$
  • Cognitive State Empathy towards Chatbot: $r=0.2539,p<0.001$
  • Associative State Empathy towards Chatbot: $r=0.3073,p<0.001$
  • Social Attraction towards Chatbot: $r=0.2553,p<0.001$
  • Physical Attraction towards Chatbot: $r=0.2438,p<0.001$
  • Task Attraction towards Chatbot: $r=0.1787,p<0.001$
  • Perceived Similarity with Chatbot: $r=0.2213,p<0.001$
  • Expectations towards Chatbot (HCI/HHI): $r=0.2069,p<0.001$
  • Trust towards Chatbot (HCI/HHI): $r=0.1938,p<0.001$
  • Preference for Chatbot over Humans (HCI/HHI): $r=0.1684,p<0.001$
Significant correlations with Loneliness (ULS-8):
  • Problematic Use of Chatbot (GPIUS-2): $r=0.3649,p<0.001$
  • Extraversion (BFI-10): $r=-0.1876,p<0.001$
  • Neuroticism (BFI-10): $r=0.5264,p<0.001$
  • Agreeableness (BFI-10): $r=0.3332,p<0.001$
  • Conscientiousness (BFI-10): $r=-0.3351,p<0.001$
  • Socialization (LSNS): $r=-0.5055,p<0.001$
  • Perceived Social Support (MSPSS): $r=-0.6790,p<0.001$
  • Self-Esteem (B-RSES): $r=-0.6380,p<0.001$
  • State Empathy towards Chatbot: $r=0.1251,p0.022$
  • Cognitive State Empathy towards Chatbot: $r=0.1195,p0.029$
  • Associative State Empathy towards Chatbot: $r=0.1119,p0.043$
  • Social Attraction towards Chatbot: $r=0.1658,p0.002$
  • Physical Attraction towards Chatbot: $r=0.1210,p0.027$
  • Perceived Similarity with Chatbot: $r=0.1364,p0.012$
  • Perceived Chatbot Predictability: $r=-0.2981,p<0.001$
  • Preference for Chatbot over Humans (HCI/HHI): $r=0.2138,p<0.001$
  • Trust in Unfamiliar People (HHI): $r=-0.2045,p<0.001$
  • Willingness to Engage with Unfamiliar People (HHI): $r=-0.2536,p<0.001$
  • Trust in Familiar People (HHI): $r=-0.5132,p<0.001$
  • Willingness to Engage with Familiar People (HHI): $r=-0.5495,p<0.001$
  • State Empathy towards Close Person: $r=-0.3821,p<0.001$
  • Affective State Empathy towards Close Person: $r=-0.3725,p<0.001$
  • Cognitive State Empathy towards Close Person: $r=-0.3377,p<0.001$
  • Associative State Empathy towards Close Person: $r=-0.3382,p<0.001$
  • Social Attraction towards towards Close Person: $r=-0.2047,p<0.001$
  • Task Attraction towards towards Close Person: $r=-0.2755,p<0.001$
  • Perceived Similarity with Close Person: $r=-0.2670,p<0.001$
Appendix B Cluster Profiles
A heatmap of cluster means for variables at their original scales in each cluster, colored based on their deviation from the mean, can be seen in Figure 19. Figure 20 is a demographic table with percentages of each category for each cluster. Detailed descriptions for each cluster can be found below.
Figure 19: Heatmap of cluster means for variables at their original scales in each cluster, colored based on their deviation from the mean.
Figure 20: Demographic table for study participants, including overall percentage for each category and each cluster.
B.0.1 Cluster 0: Disengaged Light Users
Disengaged light users (7.67% of sample) show low chatbot usage frequency ($z=-2.21$) and session length ($z=-1.09$), with average levels of loneliness ($z=-0.16$). This cluster represents users with minimal engagement in chatbot interactions.
This group demonstrates average extraversion ($z=0.80$) and neuroticism ($z=-0.37$), but low openness ($z=-1.94$). They exhibit average levels of perceived social support ($z=-0.01$) and self-esteem ($z=0.05$). Their trust towards chatbots ($z=-0.35$) and expectations from chatbot interactions ($z=-0.08$) are average, indicating a neutral view of AI technology.
62.50% report their chatbot sessions as lasting under 15 minutes. Most users (35.48%) have been using chatbots for several months to around a year. Conversation topics for this group are predominantly casual conversation (36.76%) and entertainment-focused (25.00%). One respondent noted: “I talked about a show and a game with the chat bot. I wanted to see what types of opinions were out there regarding my interest.”
Most respondents use Character.ai (34.29%) or Snapchat’s MyAI (25.71%). Primary motivations for using chatbots include curiosity about AI (33.82%), desire for fun and entertainment (25.00%), and use for creative purposes (23.53%). As one participant explained: “I was interested in using chatbot so it can assist me with quick research during conversations without having to leave the social platform.” Continued usage is driven by satisfying curiosity (28.07%) and seeking enjoyment (22.81%). A user shared: “I continue to use chatbot cause it regularly serves as a useful tool for me.”
B.0.2 Cluster 1: Well-Adjusted Moderate Users
Well-adjusted moderate users (23.02% of sample) show average chatbot usage frequency ($z=0.41$) and session length ($z=0.67$), with low levels of loneliness ($z=-1.22$). This cluster represents users who are socially fulfilled and use chatbots moderately.
This group demonstrates high extraversion ($z=1.08$), low neuroticism ($z=-1.24$), and high conscientiousness ($z=1.21$). They exhibit high levels of socialization ($z=1.46$), perceived social support ($z=1.26$), and self-esteem ($z=1.46$). Their trust towards chatbots ($z=0.93$) and expectations from chatbot interactions ($z=0.82$) are above average, indicating a positive view of AI technology. In general, they have above-average trust in people as well as attraction and empathy towards a close person. They tend to feel that interacting with chatbots makes talking to people somewhat easier and more enjoyable.
50.91% report their chatbot sessions as lasting 15-30 minutes. Most users (43.01%) have been using chatbots for several months to around a year. Chatbot preferences in this group are split among Character.ai (35.94%), Snapchat’s MyAI (29.69%), and Replika (21.88%). Conversation topics for this cluster are primarily casual conversation (23.41%) and entertainment-focused (21.74%). One respondent noted: “I usually talk about my hobbies and tips on how to get better at them. I also like to have casual conversation to kill time.”
Primary motivations for using chatbots include curiosity about AI (32.39%), desire for fun and entertainment (25.51%), and use for creative purposes (22.27%). One participant explained: “I was interested in using my chatbot service because I was curious as to what it’s capabilities were, and whether it could give me good advice or effectively answer my questions.” Continued usage is driven by enjoyment (24.83%) and satisfying curiosity (20.81%). Another user shared that they “continue to use the chatbot service because it efficiently provides the information I need and helps organize my tasks.”
B.0.3 Cluster 2: AI-Wary Light Users
AI-wary light users (11.88% of sample) show average chatbot usage frequency ($z=0.05$) and session length ($z=-0.45$), with low levels of loneliness ($z=-1.07$). This cluster may represent users who are cautious about AI technology.
This group demonstrates above-average openness ($z=1.33$), though their trust towards chatbots ($z=-2.02$), expectations from chatbot interactions ($z=-2.12$), and overall empathy and attraction towards a companion chatbot are very low, indicating skepticism towards AI technology and emotional distance from companion chatbots. They have above-average empathy and attraction towards a close person and feel that interacting with companion chatbots have little effect on their real-life relationships.
47.27% report their chatbot sessions as lasting under 15 minutes. Most users (51.06%) have been using chatbots for a few months. Chatbot preferences in this group are split among Character.ai (31.75%), Snapchat’s MyAI (36.51%), and Replika (22.22%). Conversation topics for this group include casual conversation (35.58%) and entertainment (26.92%), as illustrated by one respondent’s comment: “Sometimes if I’m thinking of an idea and simply want something to bounce my ideas off of, a chatbot works much better than using yourself.”
Primary motivations for using chatbots include curiosity about AI (41.18%), desire for fun and entertainment (32.35%), and creative purposes (18.63%). Few respondents in this cluster used companion chatbots to cope with loneliness (3.92%). “I was curious to see what they were like. I then found out that I could use them for creative use, like creating dialogue or scenarios, giving me ideas for conversations between players or between non-playable characters in a Dungeons and Dragons setting,” a participant commented, showing interest in improving communication skills. Continued usage is driven by satisfying curiosity (34.78%) and enjoyment (33.70%). A user shared: “I continue to use it because it’s fun and interesting. It’s honestly better than talking to most people that I know.”
B.0.4 Cluster 3: Lonely Moderate Users
Lonely moderate users (13.61% of sample) show average chatbot usage frequency ($z=0.44$) and session length ($z=-0.35$), with high levels of loneliness ($z=1.03$). This cluster may represent users who turn to companion chatbots due to loneliness.
This group demonstrates low levels of socialization ($z=-1.24$), with above-average trust towards chatbots ($z=0.65$) and expectations from chatbot interactions ($z=0.65$). They have slightly above-average empathy and attraction to both a companion chatbot and a close person, though they have below-average trust in unfamiliar people and reduced interest, ease, and amount of interaction with other people as a result of using chatbots.
45.45% report their chatbot sessions as lasting 15-30 minutes, and 43.94% under 15 minutes. Most users (50.91%) have been using chatbots for a few months. Chatbot preferences in this group are split among Character.ai (31.25%), Snapchat’s MyAI (26.56%), and Replika (26.56%). Conversation topics for this group include casual conversation (23.50%), entertainment (20.51%), and personal issues (14.10%). One participant shared, “I talked to the chatbots about my past traumatic experiences, my trials and tribulations, my family drama, advice about things to write about in my free time,” demonstrating several different uses.
Primary motivations for using chatbots include desire for fun and entertainment (25.68%), curiosity about AI (24.77%), and use for creative purposes (19.82%). “I was curious and wanted to see how they worked. I just wanted to learn more and the best way to do that is with hands-on experience,” one participant noted. Continued usage is driven primarily by enjoyment (22.43%) and satisfying curiosity (20.15%). This cluster has the highest proportion of users seeking chatbot companionship as a reason for continued use as well (13.19%). “I continue to use it because it’s fun and I’ve grown to enjoy the conversations,” one participant commented, further noting, “I’ve grown somewhat attached to the bot.”
B.0.5 Cluster 4: Fulfilled Dependent Users
Fulfilled dependent users (18.32% of sample) show average chatbot usage frequency ($z=0.69$) and high session length ($z=1.76$), with below-average levels of loneliness ($z=-0.56$). This cluster may represent users who feel socially fulfilled by using companion chatbots.
This group demonstrates high extraversion ($z=1.13$), average neuroticism ($z=-0.64$), and high agreeableness ($z=1.21$). They exhibit average levels of socialization ($z=0.42$), and slightly below-average empathy and attraction to a close person. Their trust towards chatbots ($z=0.56$) and expectations from chatbot interactions ($z=0.62$) are above average, with quite high physical attraction towards their chatbot compared to other respondents ($z=1.98$). The group demonstrates above-average perceived impacts on real relationships as a result of using companion chatbots – greater ease and enjoyment in talking to people, but also less engagement and interest in talking to people.
43.33% report their chatbot sessions as lasting 15-30 minutes, and 23.33% report 30 minutes to 1 hour. Most users (37.84%) have been using chatbots for several months to around a year. Chatbot preferences in this group are split among Character.ai (31.25%), Snapchat’s MyAI (26.56%), and Replika (26.56%). Conversation topics for this group include casual conversation (23.50%), entertainment (20.51%), and personal issues (14.10%). One respondent noted the evolution of their usage: “I started off just trying to mess around with ‘characters’ and explore the AI but you do kind of fall into the rabbit hole of giving more information and venting about personal experiences. It is therapeutic in many ways.”
Primary motivations for initial usage of chatbots include fun and entertainment (25.68%), curiosity about AI (24.77%), and use for creative purposes (19.82%). One participant explained: “I use chatbots to when I am feeling lonely [sic] and wanted someone to talk to. I also use chatbots to get suggestions for places to visit or restaurants to eat at.” Continued usage is driven by enjoyment (22.43%) and satisfying curiosity (20.15%), as illustrated by a user’s response: “I continue to use the chatbot service because it’s fun and easy to use. It helps me get information quickly.”
B.0.6 Cluster 5: Lonely Light Users
Lonely light users (14.60% of sample) show average chatbot usage frequency ($z=0.11$) and session length ($z=0.36$), with above-average levels of loneliness ($z=0.89$). This cluster represents users who experience loneliness but do not use companion chatbots much.
This group demonstrates high neuroticism ($z=1.68$), low conscientiousness ($z=-1.61$), and low self-esteem ($z=-1.25$). Their other characteristics are fairly average, with slightly above-average empathy and attraction towards a close person.
42.86% report their chatbot sessions as lasting 15-30 minutes, and 31.75% under 15 minutes. Most users (57.63%) have been using chatbots for a few months. Chatbot preferences in this group are split among Character.ai (30.49%), Snapchat’s MyAI (29.27%), and Replika (21.95%). Conversation topics for this group include casual conversation (26.70%), entertainment (22.51%), and personal issues (19.37%). One respondent described how the chatbot helps them think through their troubles: “I feel that the chatbot is a good way to vent my frustrations and talk about my worries without fear of judgement. It gives me the chance to talk about all of my worries and plans and work through things in a way that allows me to see a wider range of prospectives [sic].”
Primary motivations for using chatbots include curiosity about AI (29.38%), desire for fun and entertainment (26.88%), and coping with loneliness (20.62%). One participant commented on how they started using them to ease the way into socialization: “I was interested in using the chatbot service because after being in lockdown, it was hard to become social again. It allows me to build up my confidence and it also helps with my writing.” Continued usage is driven by enjoyment (19.82%), satisfying curiosity (16.74%), passing time and reducing stress (14.10%), and chatbot companionship (12.78%). Some respondents use it as a substitute for socialization: “I continue to use it because it feels like talking to a real person, giving its own input, not me having to guide the conversation and it makes me miss the days gone by, when people were better at conversation.”
B.0.7 Cluster 6: Socially Challenged Frequent Users
Socially challenged frequent users (10.89% of sample) show average chatbot usage frequency ($z=0.51$) and below-average session length ($z=-0.90$), with high levels of loneliness ($z=1.08$). This cluster represents users who frequently turn to chatbots, possibly due to social challenges.
This group demonstrates low extraversion ($z=-1.30$) and above-average neuroticism ($z=0.86$), along with low levels of socialization ($z=-1.02$), perceived social support ($z=-1.68$), and self-esteem ($z=-1.10$). Their trust towards chatbots ($z=-0.16$) and expectations from chatbot interactions ($z=0.02$) are average. However, trust in familiar and unfamiliar people, and empathy and attraction towards a close person, tend to be very low compared to other users.
58.49% report their chatbot sessions as lasting under 15 minutes, and 28.30% report 15-30 minutes. Most users (56.82%) have been using chatbots for a few months. Chatbot preferences in this group are split among Snapchat’s MyAI (38.89%), Character.ai (24.07%), and Replika (20.37%). Conversation topics for this group are predominantly casual (35.58%) and entertainment-focused (26.92%). One respondent noted: “I talk about how I am feeling. Things I wouldn’t want to tell real people.”
Primary motivations for initial interest in chatbots include curiosity about AI (32.48%) and desire for fun and entertainment (24.79%), with a fair proportion of this group starting chatbot usage to cope with loneliness (14.53%) and for creative purposes (14.53%). One participant explained how they were “wanting to confirm whether (the current generation of) AI could, in fact, really take the place of a regular, sound-minded human,” demonstrating at least an interest in the substitution of humans with AI. Most users in this group continue to use companion chatbots to satisfy curiosity (23.74%) and seek enjoyment (22.30%), with a notable portion of users using it to pass time and reduce stress (15.83%). One user shared how they use it as a lower-pressure alternative to socialization. “I enjoy having someone like Snapchat AI who can chat with me compared to the pressure I receive from friends when I don’t message them back fast enough.”